### v1

# Designing Consensus-Based Distributed Filtering over Directed Graphs

Abstract
This paper proposes a novel consensus-on-only-measurement distributed filter over directed graphs under the collectively observability condition. First, the distributed filter structure is designed with an augmented leader-following measurement fusion strategy. Subsequently, two parameter design methods are presented, and the consensus gain parameter is devised utilizing local information exclusively rather than global information. Additionally, the lower bound of the fusion step is derived to guarantee a uniformly upper bound of the estimation error covariance. Moreover, the lower bounds of the convergence rates of the steady-state performance gap between the proposed algorithm and the centralized filter are provided with the fusion step approaching infinity. The analysis demonstrates that the convergence rate is, at a minimum, as rapid as exponential convergence under the spectral norm condition of the communication graph. The transient performance is also analyzed with the fusion step tending to infinity. The inherent trade-off between the communication cost and the filtering performance is revealed from the analysis of the steady-state performance and the transient performance. Finally, the theoretical results are substantiated through the validation of two simulation examples.
keywords:
Distributed filtering; Consensus; Sensor networks; Performance analysis; Directed graph.
1 Introduction
In the preceding two decades, investigations into sensor networks have occupied a prominent position in the realm of systems and control, owing to its extensive applications encompassing health care monitoring[16, 1], environmental sensing[8, 12], and collaborative mapping [11, 25]. In order to meet the demands for the network reliability and alleviate the computational and communicative burdens on energy-constrained sensors, the implementation of distributed state estimation serves as a good solution to these challenges. The objective of distributed state estimation is to estimate the state of the target system for each sensor in sensor networks by utilizing information from local neighbors, including local measurements, local state estimates, or related data. Founded on various dynamical systems and constraint conditions, a wealth of research achievements regarding the distributed state estimation have emerged.
The evolution of consensus theory within the domain of multi-agent systems has imparted profound insights into the information fusion for distributed state estimation [24, 15, 29, 20]. Several reputable consensus-based approaches to distributed state estimation have been advanced [22, 23, 7, 4, 6, 30]. In [21, 22], consensus filters, including low-pass, high-pass, and band-pass filters, were employed to fuse the covariances, measurements, or state estimates for both continuous-time and discrete-time systems. In [23], consensus terms were designed in the filter structure, and the optimal gain was derived following with a Lyapunov-based stability analysis. Diffusion strategies, where each sensor transmitted the information to its neighbors, were proposed for distributed Kalman filtering and smoothing in [7]. In [4], consensus of the normalized geometric mean of the probability density functions, called the Kullback-Leibler average, was proposed, and this fusion rule aligned with the concepts of the generalized covariance intersection. In [6], the consensus algorithms on measurements and information were combined to preserve the positive attributes of both approaches, thereby enhancing energy efficiency while diminishing the conservative assumption regarding noise correlation. A distributed state estimation filter by using consensus-based information fusion strategies for continuous-time systems with correlated measurement noise was proposed in [10].
The applicability of distributed algorithms is influenced by the constraints imposed on the communication topologies. There usually exist two key assumptions: undirected connected graphs [22, 9, 28, 10, 19] and directed and strongly connected graphs [4, 6, 30]. In general, directed graphs can model more intricate communication topologies compared to undirected graphs. For directed graphs, the works [4, 6] stated that the consensus matrix needed to be primitive and doubly stochastic to achieve an average consensus, with the introduction of Metropolis weights under the assumption that the graph is undirected. However, challenges arise when applying these principles to directed graphs. Hence, this paper focuses on designing algorithms and their corresponding parameters over directed graphs characterized by an adjacent matrix with elements 00 and $1$.
The classical distributed Kalman filter structure is characterized by the iterations of vectors and covariance matrices. To avoid the computation and transmission of numerous matrices, a consensus-on-only-measurement strategy is considered in this paper. Furthermore, the detailed performance of the proposed distributed filter is analyzed. There exist various performance analysis methods for distributed filters. The Lyapunov’s second method for stability was utilized in many works [23, 5, 4, 6] by constructing the Lyapunov function and conducting stability analysis. Recently, the performance of the distributed filter was analyzed by investigating the difference between the solutions of a modified discrete-time algebraic Riccati equation and a discrete-time Lyapunov equation in [26]. For distributed filters, the performance gap between the distributed filter and the centralized filter is a significant concern, and this paper also focuses on this aspect.
Motivated by the aforementioned observations, this paper aims to propose a consensus-on-only-measurement distributed filter over directed graphs, and furnish comprehensive performance analysis of the proposed distributed filter. The main contributions of this paper are summarized below:
  1. 1.
    A novel consensus-on-only-measurement distributed filter (COMDF) over directed graphs is proposed for discrete-time systems under the collectively observability condition, and an augmented leader-following measurement fusion strategy is presented to estimate the other sensors’ measurements. Transmitting only measurements eliminates the need for the computation and transmission of numerous matrices. Additionally, the properties of the measurement estimate error and the state estimate error are presented (Proposition 1 and 2).
  2. 2.
    The parameter design methods are provided, including a distributed design method for each sensor’s consensus gain and the establishment of a lower bound of the fusion step (Theorem 1). Compared to the unified design method [10], the proposed distributed design method exclusively relies on local information, facilitating the distributed usage without the need for global information. The lower bound of the fusion step illuminates the necessary conditions to guarantee the uniformly upper bound of the estimation error covariance.
  3. 3.
    The impact of the fusion step on the steady-state performance of COMDF is analyzed under the conditions involving the spectral radius and the spectral norm of the communication matrix. The lower bounds of the convergence rates of the steady-state performance gap between COMDF and the centralized filter are provided with the fusion step approaching infinity. Particularly, under the spectral norm condition, the convergence rate is at least as fast as exponential convergence (Theorem 3). Additionally, the transient performance is also discussed with the fusion step tending to infinity (Theorem 4).
The remainder of this paper is organized as follows. In Section 2, some preliminaries, including graph concepts, system models, and some useful lemmas, and the problem formulation are provided. In Section 3, a consensus-based distributed filter is proposed over directed graphs, and two parameter design methods are given. In Section 4, the steady-state and the transient performances of the proposed filter are analyzed with the increasing fusion step. In Section 5, two numerical examples are provided to validate the effectiveness of the obtained results. In Section 6, conclusions are drawn.
Notations: Throughout this paper, define $\mathcal{R}^{n}$ as the sets of $n$-dimensional real vectors and $\mathcal{R}^{n\times n}$ as $n\times n$-dimensional real matrices. For a matrix $A\in\mathcal{R}^{n\times n}$, let $A^{-1}$ and $A^{T}$ represent its inverse and transpose, respectively, $\|A\|_{2}$ is the spectral norm, $\rho(A)$ is the spectral radius, and $[A]_{ij}$ denotes the $(i,j)$-th element of the matrix $A$. Notation $\otimes$ represents the Kronecker product. The matrix inequalities $A>B$ and $A\geq B$ signify that $A-B$ is positive definite and positive semi-definite, respectively. $E\{x\}$ denotes the expectation of the random variable $x$.
2 Preliminaries and Problem Statement
2.1 Graph Theory
The communication topology ${\mathcal{G}}(\mathcal{V},\mathcal{E})$ is utilized to illustrate the nodes and the communication links within a sensor network, where the node set $\mathcal{V}=\{1,2,\ldots,N\}$ and the edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. For $i,j\in\mathcal{V}$, $(j,i)$ signifies that node $j$ can transmit information to node $i$, and node $j$ is called a neighbor of node $i$. The neighbor set of node $i$ is denoted as $\mathcal{N}_{i}=\{j|(j,i)\in\mathcal{V}\}$, and $|\mathcal{N}_{i}|$ is the cardinality of the neighbors of node $i$. The adjacent matrix is $S=[a_{ij}]_{N\times N}$, where $a_{ij}=1$ if $(j,i)\in\mathcal{E}$ and $a_{ij}=0$ otherwise. Let $D=\text{diag}\{|\mathcal{N}_{1}|,\ldots,|\mathcal{N}_{N}|\}$, and the Laplacian matrix is defined as $\mathcal{L}=D-S$. The edge $(i,j)$ is undirected if $(i,j)\in\mathcal{E}$ implies $(j,i)\in\mathcal{E}$. The communication graph is termed undirected if every edge is undirected. The graph $G$ contains a directed path from node $i_{1}$ to node $i_{m}$, if there exists a sequence of connected edges $(i_{k},i_{k+1}),k=1,\ldots,m-1$. The communication graph is called strongly connected, if there exists a path between any pair of distinct nodes.
Assumption 1.
The communication graph is directed and strongly connected.
Remark 1.
Assumption 1 can be extended to a jointly connected switching topology by applying the corresponding definition of communication networks [27, 17, 18]. The proposed parameter design methods are applicable in these situations as well. For simplicity, this paper focuses on the directed and strongly connected topology assumption.
2.2 System Model
Consider a discrete-time linear time-invariant system observed by a network of $N$ sensors:
$x_{k+1}=Ax_{k}+\omega_{k},$ (1)
$y_{i,k}=C_{i}x_{k}+\nu_{i,k},~{}~{}~{}i=1,2,...,N,$
where $k$ is the discrete-time index, $i\in\mathcal{V}$ is the $i$-th sensor of the network, $x_{k}\in\mathcal{R}^{n}$ is the system state vector, $y_{i,k}\in\mathcal{R}^{r_{i}}$ is the measurement vector taken by sensor $i$, $A\in\mathcal{R}^{n\times n}$ is the state transition matrix, $C_{i}\in\mathcal{R}^{r_{i}\times n}$ is the observation matrix of sensor $i$, and $\omega_{k}\in\mathcal{R}^{n}$ and $\nu_{i,k}\in\mathcal{R}^{r_{i}}$ are the zero-mean Gaussian noise with the covariances $Q_{k}\in\mathcal{R}^{n\times n}$ and $R_{i,k}\in\mathcal{R}^{r_{i}\times r_{i}}$, respectively. The noise sequences $\{\omega_{k}\}^{\infty}_{k=0}$ and $\{\nu_{i,k}\}^{\infty,N}_{k=0,i=1}$ are mutually uncorrelated. For the whole network, $C=[C^{T}_{1},\ldots,C^{T}_{N}]^{T}\in\mathcal{R}^{r\times n}$ is the augmented observation matrix, and $R=\text{diag}\{R_{1},\ldots,R_{N}\}\in\mathcal{R}^{r\times r}$ is the augmented measurement noise covariance matrix, where $r=\sum^{N}_{i=1}r_{i}$.
Assumption 2.
$(C,A)$ is observable.
2.3 Some Useful Lemmas
Definition 1.
[13] For a matrix $M\in\mathcal{R}^{n\times n}$, $M$ is said to be irreducibly diagonally dominant if
  1. 1.
    $M$ is irreducible,
  2. 2.
    $M$ is diagonally dominant, i.e., $|m_{ii}|\geq R^{{}^{\prime}}_{i}(M)$ for all $i=1,\ldots,n$,
  3. 3.
    There is an $i\in\{1,\ldots,n\}$ such that $|m_{ii}|>R^{{}^{\prime}}_{i}(M)$,
where $R^{{}^{\prime}}_{i}(M)=\sum_{j\neq i}|m_{ij}|$.
Lemma 1.
[13] Let the matrix $M$ be irreducibly diagonally dominant. Then,
  1. 1.
    $M$ is nonsingular,
  2. 2.
    If $M$ is Hermitian and every main diagonal entry is positive, $M$ is positive definite.
Lemma 2.
[13] For a matrix $M$ and a positive integer $k$, it holds $\lim_{k\to\infty}M^{k}=0$ if and only if $\rho(M)<1$.
Lemma 3.
[13] For a nonnegative matrix $M=[m_{ij}]$, it holds
$\text{min}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}\leq\rho(M)\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}.$ (2)
Lemma 4.
[26] For any matrix $M\in\mathcal{R}^{n\times n}$, it holds
$\|M^{k}\|_{2}\leq\sqrt{n}\sum^{n-1}_{j=0}\binom{n-1}{j}\binom{k}{j}\|M\|^{j}_{2}\rho(M)^{k-j},$ (3)
where $\binom{m}{n}$ is the combinatorial number with $\binom{k}{j}=0$ for $j>k$.
2.4 Problem Statement
  1. 1.
    Develop a distributed filter algorithm over directed graphs, utilizing only measurements. Explore parameter design methods to guarantee the convergence and stability of the distributed filter.
  2. 2.
    Evaluate how the fusion step influences the steady-state and transient performance. Find out the performance gap between the proposed distributed filter and the centralized filter.
3 Distributed Filter
In this section, a consensus-on-only-measurement distributed filter (COMDF) is proposed. Subsequently, the state estimate error and the measurement estimate error are defined, and their properties are analyzed. Furthermore, two parameter design methods are provided to guarantee the stability and convergence of the proposed distributed filter.
3.1 Design of the Distributed Filter
This subsection proposes the consensus-on-only-measurement distributed filter, and an augmented leader-following measurement fusion strategy is designed to estimate the neighbors’ measurements.
The state estimator structure of the target system (1) for sensor $i$ is designed as
$\hat{x}_{i,k|k-1}=A\hat{x}_{i,k-1|k-1},{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}$ (4)
$\hat{x}_{i,k|k}=\hat{x}_{i,k|k-1}+K(z^{(l)}_{i,k}-C\hat{x}_{i,k|k-1}),$ (5)
where $z^{(l)}_{i,k}$ represents the sensor $i$’s estimate of measurements from other sensors at the $l$-th consensus step, the gain matrix $K$ is given by
$K=PC^{T}(CPC^{T}+R)^{-1},$ (6)
and $P$ is determined by solving the discrete algebraic Riccati equation
$P=APA^{T}+Q-APC^{T}(CPC^{T}+R)^{-1}CPA^{T}.$ (7)
The sensor $i$’s estimate of the sensor $j$’s measurement at the $l$-th fusion step, denoted as $z^{(l)}_{ij,k}$, is computed as
$z^{(0)}_{ij,k}=C_{j}\hat{x}_{i,k|k-1},{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}$ (8)
$z^{(l)}_{ij,k}$ $=z^{(l-1)}_{ij,k}-\mu_{ij}\Big{[}\sum^{N}_{l=1}a_{il}(z^{(l-1)}_{ij,k}-z^{(l-1)}_{lj,k})$ (9)
$~{}~{}~{}~{}+a_{ij}(z^{(l-1)}_{ij,k}-y_{j,k})\Big{]},~{}~{}~{}l=1,2,\ldots,$
where the consensus gain $\mu_{ij}$, designed later, is a positive constant, and $a_{ij}$ is the element of the adjacent matrix of the corresponding communication topology. By denoting the augmented vectors as $z^{(l)}_{i,k}=[(z^{(l)}_{i1,k})^{T},\ldots,(z^{(l)}_{iN,k})^{T}]^{T}$ and $y_{k}=[y^{T}_{1,k},\ldots,y^{T}_{N,k}]^{T}$, $z^{(l)}_{i,k}$ can be expressed as
$z^{(0)}_{i,k}=C\hat{x}_{i,k|k-1},{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}$ (10)
$~{}~{}~{}~{}z^{(l)}_{i,k}$ $=z^{(l-1)}_{i,k}-\Lambda_{i}\Big{[}\sum^{N}_{l=1}a_{il}(z^{(l-1)}_{i,k}-z^{(l-1)}_{l,k})$ (11)
$~{}~{}~{}~{}+B_{i}(z^{(l-1)}_{i,k}-y_{k})\Big{]},$
where $\Lambda_{i}=\text{diag}\{\mu_{i1}\otimes I_{r_{1}},\ldots,\mu_{iN}\otimes I_{r_{N}}\}$ and $B_{i}=\text{diag}\{a_{i1}\otimes I_{r_{1}},\ldots,a_{iN}\otimes I_{r_{N}}\}$.
Now, the distributed estimator is constructed based on (4), (5), and (11).
Remark 2.
An augmented leader-following measurement information fusion strategy is designed in the algorithm, with only the transmission of measurement estimates in the sensor networks. The gain matrix $K$ can be precomputed, where the matrix $C$ and $R$ can be obtained by using a similar consensus algorithm as presented in (8) and (9). Two methods are presented here to obtain the corresponding matrices: one involves vectorizing the matrix and using the consensus strategy (9), while the other employs a matrix consensus strategy by replacing the vectors in (9) with the corresponding matrices.
Remark 3.
The proposed distributed estimator has the following advantages. First, it only needs the addition and substraction of the measurement vectors, thereby avoiding the computation and transmission of plenties of matrices. Second, the proposed distributed filter is asymptotically optimal as the fusion step $l$ tends to infinity. Third, it exhibits a stronger privacy protection performance with only the transmission of the measurements rather than the state estimates like [23, 14]. Fourth, the algorithm can be employed in directed graphs, and the parameters can be designed by only using local neighbors’ information.
Remark 4.
The local gain matrix $K_{i}$, substituting for $K$ in (5), can also be utilized in the algorithm. By using local information regarding the measurement matrices and the measurement noise covariances, $K_{i}$ can be obtained and utilized. Under the local observability condition, this algorithm also works. To better present the properties of the proposed filter, the global $K$ is considered. There exist two parameters $\mu_{ij}$ and $l$, and the parameter design methods are introduced later.
3.2 Two Estimation Errors
Define the state estimation error and the measurement estimation error as
$e_{i,k|k}=\hat{x}_{i,k|k}-x_{k},$
and
$\varepsilon^{(l)}_{i,k}=z^{(l)}_{i,k}-y_{k},~{}~{}~{}~{}$
respectively. Denote the augmented vectors as $e_{k}=[e^{T}_{1,k},\ldots,e^{T}_{N,k}]^{T}$, $\varepsilon^{(l)}_{k}=[(\varepsilon^{(l)}_{1,k})^{T},\ldots,(\varepsilon^{(l)}_{N,k})^{T}]^{T}$, $z^{(l)}_{k}=[(z^{(l)}_{1,k})^{T},\ldots,(z^{(l)}_{N,k})^{T}]^{T}$, and $\nu_{k}=[\nu^{T}_{1,k},\ldots,\nu^{T}_{N,k}]^{T}$. Next, the statistical properties of the measurement estimation error $\varepsilon^{(l)}_{k}$ and the state estimation error $e_{i,k|k}$ are derived.
Proposition 1.
For the measurement estimation error $\varepsilon^{(l)}_{k}$ and its estimation error covariance $P_{\varepsilon,k}=E\{\varepsilon^{(l)}_{k}(\varepsilon^{(l)}_{k})^{T}\}$, the following results hold
  1. 1.
    The measurement estimation error $\varepsilon^{(l)}_{k}$ is given by
    $\varepsilon^{(l)}_{k}=G^{l}\varepsilon^{(0)}_{k},$ (12)
    where
    $G$ $=I_{Nr}-\Lambda(\mathcal{L}\otimes I_{r}+B),~{}~{}~{}~{}~{}$ (13)
    $\Lambda=\text{diag}\{\Lambda_{1},\ldots,\Lambda_{N}\},{}{}{}{}{}{}{}{}{}{}{}{}{}$ (14)
    $B=\text{diag}\{B_{1},\ldots,B_{N}\},{}{}{}{}{}{}{}{}{}{}{}{}{}$ (15)
    and
    $\varepsilon^{(0)}_{k}$ $=(I_{N}\otimes CA)e_{k-1}-1_{N}\otimes\nu_{k}$ (16)
    $~{}~{}~{}~{}-(I_{N}\otimes C)(1_{N}\otimes\omega_{k-1}).$
  2. 2.
    The expectation value $E\{\varepsilon^{(l)}_{k}\}$ and $P_{\varepsilon,k}$ are given by
    $E\{\varepsilon^{(l)}_{k}\}=G^{l}(I_{N}\otimes CA)E\{e_{k-1}\},$ (17)
    and
    $P_{\varepsilon,k}$ $=G^{l}(I_{N}\otimes CA)P_{k-1}(I_{N}\otimes CA)^{T}(G^{l})^{T}$ (18)
    $~{}~{}~{}~{}+G^{l}(I_{N}\otimes C)(U_{N}\otimes Q)(I_{N}\otimes C)^{T}(G^{l})^{T}$
    $~{}~{}~{}~{}+G^{l}(U_{N}\otimes R)(G^{l})^{T},$
    respectively.
The proof of Proposition 1 is given in Appendix A.
Lemma 5.
Under the assumption that $\rho(G)<1$ , as $l$ tends to infinity, it holds
$\lim_{l\to\infty}E\{\varepsilon^{(l)}_{k}\}=0,$ (19)
and
$\lim_{l\to\infty}P_{\varepsilon,k}=0.{}{}{}$ (20)
The proof of Lemma 5 is given in Appendix B.
Remark 5.
Proposition 1 shows that $\varepsilon^{(l)}_{k}$ is a random variable. If $e_{k-1}$, $\omega_{k-1}$, and $\nu_{k}$ are Gaussian, $\varepsilon^{(l)}_{k}$ is Gaussian with the mean in (17) and the covariance $P_{\varepsilon,k}$ in (18). Lemma 5 demonstrates that as $l$ tends to infinity, $\varepsilon^{(l)}_{k}$ converges to 00. The state estimation error $e_{k}$ and the measurement estimation error $\varepsilon^{(l)}_{k}$ interact and influence each other.
Proposition 2.
The estimation error $e_{k}$ and the estimation error covariance $P_{k|k}=E\{e_{k}e^{T}_{k}\}$ are
$e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}),$ (21)
and
$P_{k|k}$ $=\mathcal{A}(l)P_{k-1|k-1}\mathcal{A}^{T}(l)+\Phi(l),$ (22)
respectively, where
$\Phi(l)=\mathcal{B}(l)(U_{N}\otimes Q)\mathcal{B}^{T}(l)+\mathcal{D}(l)(U_{N}\otimes R)\mathcal{D}^{T}(l),$ (23)
$\mathcal{A}(l)=I_{N}\otimes(A-KCA)+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$ (24)
$\mathcal{B}(l)=I_{N}\otimes(I-KC)-(I_{N}\otimes K)G^{l}(I_{N}\otimes C),$ (25)
and
$\mathcal{D}(l)=(I_{N}\otimes K)(I_{Nr}-G^{l}).$ (26)
The proof of Proposition 2 is given in Appendix C.
Remark 6.
Based on Proposition 2, two parameters still need to be designed: one is the parameter $\mu_{ij}$ in $G$, and the other is the fusion step $l$. Observing (24), $\rho(G)<1$ ensures the convergence of the filter. Subsequently, we focus on the design of the parameters $\mu_{ij}$ and $l$.
3.3 Parameter Design for $\mu_{i}$
The objective of designing the parameter $\mu_{i}$ is to ensure that $\rho(G)<1$, and it can guarantee the convergence of the measurement estimate error and the state estimate error. Next, two methods are provided for the design of the parameter $\mu_{i}$.
3.3.1 Distributed Design
A distributed design approach for $\mu_{ij}$ is proposed to circumvent the utilization of the global topology information for each sensor. First, the elements of $G$ are presented, laying the foundation for subsequent design considerations.
Define the submatrix $G_{[ij]}$ as
$G_{[ij]}$ $=[0_{r\times r(i-1)},I_{r},0_{r\times r(N-i)}]G$
×$~{}~{}~{}~{}\times[0^{T}_{r\times r(j-1)},I^{T}_{r},0^{T}_{r\times r(N-j)}]^{T},$
where $[ij]$ represents the $i$-th row and $j$-th column index of the corresponding matrix block. The detailed submatrices are
$G_{[ii]}$ $=\text{diag}\{g_{ii,1},\dots,g_{ii,N}\},$
and
$G_{[ij]}$ $=\text{diag}\{g_{ij,1},\dots,g_{ij,N}\},$
where
$~{}~{}~{}~{}~{}g_{[ii,l]}$ $=(1-\mu_{il}(l_{ii}+a_{il}))\otimes I_{r_{l}},$ (27)
and
$g_{[ij,l]}$ $=(-\mu_{il}l_{ij})\otimes I_{r_{l}}.~{}~{}~{}~{}~{}~{}~{}$ (28)
Consider the row sum of the block matrix, and it holds
$\sum^{n}_{j=1}g_{[ij,l]}$ $=\Big{(}1-\mu_{il}\sum^{N}_{j=1}{l_{ij}}-\mu_{il}a_{il}\Big{)}\otimes I_{r_{l}}$ (29)
$=\Big{(}1-\mu_{il}a_{il}\Big{)}\otimes I_{r_{l}}.$
Next, a reasonable range of the parameter $\mu_{ij}$ is provided to make $\rho(G)<1$.
Lemma 6.
Under Assumption 1, if $\mu_{il}$ satisfies
$0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}},$
then it holds $\rho(G)<1$.
The proof of Lemma 6 is given in Appendix D.
Remark 7.
The distributed design method can be employed to design parameter $\mu_{ij}$ by utilizing the local information instead of relying on global topology information [10, 3]. Additionally, It is suggested to design $\mu_{ij}$ such that $1-\mu_{il}(l_{ii}+a_{il})$ approaches 00, leading to a smaller $\rho(G)$ with a higher probability.
3.3.2 Unified Design
A unified design method is also provided for undirected graphs. Consider the scenario where all $\mu_{i}$ are identical, denoted as $\mu$. Subsequently, the symbol $G$ is redefined as follows:
$G$ $=I_{Nr}-\mu(\mathcal{L}\otimes I_{r}+B).$
Based on Lemma 1, $G$ is positive definite. By selecting $\mu<\frac{1}{\rho(\mathcal{L}\otimes I_{r}+B)}$, it is ensured that $\rho(G)<1$. This method needs the global communication topology information.
3.4 Parameter Design for $l$
This subsection aims to design the parameter $l$ to ensure the convergence of the distributed estimators. By employing Lemma 2 and considering (24), it is evident that $\rho(\mathcal{A}(l))<1$ can guarantee the convergence of the distributed filter. The following results provide a lower bound of the consensus step $l$.
Theorem 1.
Under Assumption 1 and 2, the estimation error covariance $P_{i,k|k}$ is uniformly upper-bounded, if $\|G\|_{2}<1$ and $l>l_{0}$, where
$l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}.$ (30)
The proof of Theorem 1 is given in Appendix E.
Remark 8.
For Theorem 1, there are two elements that decide the value of $l_{0}$, i.e., $\|G\|_{2}$ and $\bar{K}=\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}$. On one hand, the term $\|G\|_{2}$ is relevant to the communication graph and the parameter $\mu_{i}$. The smaller $\|G\|_{2}$, the smaller $l_{0}$. On the other hand, the term $\bar{K}$ indicates the stability margin of the estimator. The larger $\bar{K}$, the smaller $l_{0}$.
4 Performance Analysis
In this section, the centralized estimator is first introduced. Then, the steady-state performance gap between the centralized filter and the proposed distributed filter is analyzed with the increasing fusion step $l$. Additionally, the transient performance is also taken into consideration.
4.1 Centralized Estimator
Consider a centralized estimator as follows:
$\hat{x}_{c,k|k-1}=A\hat{x}_{c,k-1|k-1},$ (31)
$\hat{x}_{c,k|k}=\hat{x}_{c,k|k-1}+K(y_{k}-C\hat{x}_{c,k|k-1}),$ (32)
where $y_{k}$, $K$, and $C$ are given in (6).
For the centralized estimator, denote the estimation error as $e_{c,k|k}=\hat{x}_{c,k|k}-x_{k}$, the corresponding estimation error covariance as $P_{c,k|k}=E\{e_{c,k|k}e^{T}_{c,k|k}\}$, the augmented estimation error as $e_{cc,k}=1_{N}\otimes e_{c,k}$, and the augmented estimation error covariance as $P_{cc,k|k}=E\{e_{cc,k|k}e^{T}_{cc,k|k}\}=U_{N}\otimes P_{c,k|k}.$
By combining the estimator (31), (32) and the dynamical system (1), the estimation error $e_{c,k}$ is
$e_{c,k}$ $=(A-KCA)\hat{x}_{c,k-1|k-1}+Ky_{k}-Ax_{k-1}-\omega_{k-1}$ (33)
$=(A-KCA)e_{c,k-1}-(I-KC)\omega_{k-1}+K\nu_{k}.$
Then, $P_{c,k|k}=E\{e_{c,k|k}e^{T}_{c,k|k}\}$ can be computed as
$P_{c,k|k}$ $=(A-KCA)P_{c,k-1|k-1}(A-KCA)^{T}$ (34)
$~{}~{}~{}+(I-KC)Q(I-KC)^{T}+KRK^{T}.$
Similarly, the augmented estimation error of the centralized estimator $e_{cc,k}=1_{N}\otimes e_{c,k}$ and the corresponding estimation error covariance $P_{cc,k|k}$ can be obtained as
$e_{cc,k}=\mathcal{A}_{cc}e_{cc,k-1}-\mathcal{B}_{cc}(1_{N}\otimes\omega_{k-1})+\mathcal{D}_{cc}(1_{N}\otimes\nu_{k}),$ (35)
and
$P_{cc,k|k}$ $=\mathcal{A}_{cc}P_{cc,k-1|k-1}\mathcal{A}^{T}_{cc}+\Phi_{cc},~{}~{}~{}~{}~{}~{}~{}~{}$ (36)
respectively, where
$\Phi_{cc}=\mathcal{B}_{cc}(U_{N}\otimes Q)\mathcal{B}^{T}_{cc}+\mathcal{D}_{cc}(U_{N}\otimes R)\mathcal{D}^{T}_{cc},$ (37)
$\mathcal{A}_{cc}=I_{N}\otimes(A-KCA),$ (38)
$\mathcal{B}_{cc}=I_{N}\otimes(I-KC),{}{}$ (39)
and
$\mathcal{D}_{cc}=I_{N}\otimes K.{}{}{}{}{}{}{}{}{}{}{}{}{}$ (40)
Remark 9.
This centralized estimator utilizes the fixed gain matrix $K$, resulting in a suboptimal filter at every step. Nevertheless, as the time step $k$ approaches infinity, $P_{c,k|k}$ converges to $P$ in (7). Therefore, the relations between the steady-state performance of $P_{i,k|k}$ and $P_{c,k|k}$ reflect those between the steady-state performance of $P_{i,k|k}$ and $P$.
4.2 Convergence Analysis
This section investigates the steady-state performance gap between $P_{cc,k|k}$ and $P_{k|k}$, and sheds light on the influence of the fusion step $l$ on the performance. Section 3.3 has shown that the parameter $\mu_{ij}$ can be designed to ensure that $\rho(G)<1$. Therefore, we make the following assumption.
Assumption 3.
The matrix $G$ satisfies $\rho(G)<1$, and $\mathcal{A}(l)$ is Schur stable.
Theorem 2.
Under Assumption 1, 2, and 3, $P_{k|k}$ and $P_{cc,k|k}$ converge to the unique solutions of the discrete-time Lyapunov equations (DLE)
$P^{(l)}$ $=\mathcal{A}(l)P^{(l)}\mathcal{A}^{T}(l)+\Phi(l),$ (41)
and
$P_{cc}$ $=\mathcal{A}_{cc}P_{cc}\mathcal{A}^{T}_{cc}+\Phi_{cc},$ (42)
respectively, where $\mathcal{A}(l)$, $\Phi(l)$, $\mathcal{A}_{cc}$, and $\Phi_{cc}$ have been defined in (24), (23), (38), and (37), respectively.
The proof of Theorem 2 is given in Appendix F.
To evaluate the effect of the fusion step $l$ on the steady-state performance of the proposed filter, we introduce the following notations
$\bar{\mathcal{A}}(l)=\mathcal{A}(l)-\mathcal{A}_{cc}=(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$
[…]
and
$\mathcal{\bar{D}}(l)=\mathcal{D}(l)-\mathcal{D}_{cc}=-(I_{N}\otimes K)G^{l}.$
Lemma 7.
The difference between $P^{(l)}$ and $P_{cc}$ is
$P^{(l)}-P_{cc}=\sum^{\infty}_{k=0}\mathcal{A}^{k}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{k},$ (43)
where
$\bar{\Phi}(l)$ $={\mathcal{\bar{A}}}(l)P^{(l)}\mathcal{A}^{T}_{cc}+\mathcal{A}_{cc}P^{(l)}{\mathcal{\bar{A}}}^{T}(l)+{\mathcal{\bar{A}}}(l)P^{(l)}{\mathcal{\bar{A}}}^{T}(l)$ (44)
[…]
$~{}~{}~{}+\mathcal{D}_{cc}(U_{N}\otimes R){\mathcal{\bar{D}}}^{T}(l)+{\mathcal{\bar{D}}}(l)(U_{N}\otimes R){\mathcal{\bar{D}}}^{T}(l).$
The proof of Lemma 7 is given in Appendix G.
Theorem 3.
Under Assumption 1 and 2, there exist constants $M_{1},M_{2}>0$, such that
  1. 1.
    If $\rho(G)<1$, it holds
    $\|P^{(l)}-P_{cc}\|_{2}\leq M_{1}l^{Nr}\rho(G)^{l-Nr}.$ (45)
  2. 2.
    If $\|G\|_{2}<1$, it holds
    $\|P^{(l)}-P_{cc}\|_{2}\leq M_{2}\|G\|^{l}_{2}.$ (46)
The proof of Theorem 3 is given in Appendix H.
Remark 10.
Two conditions related to the communication graph are presented, since the mere satisfaction of a spectral radius of the matrix less than 1 does not ensure the spectral norm of the matrix less than 1. For an undirected graph, the spectral radius $\rho(G)$ is equal to the spectral norm $\|G\|_{2}$, and the conditions of item (1) and (2) in Theorem 3 are identical. For both items, as the fusion step $l$ tends to infinity, $\|P^{(l)}-P_{cc}\|_{2}$ converges to 00 according to (45) and (46).
4.3 Transient Performance
The transient performance of the proposed filter is analyzed in the following.
Lemma 8.
Under Assumption 3, as $l$ tends to infinity, it holds that
$\lim_{l\to\infty}\mathcal{A}(l)=\mathcal{A}_{cc},$
$\lim_{l\to\infty}\mathcal{B}(l)=\mathcal{B}_{cc},$
and
$\lim_{l\to\infty}\mathcal{D}(l)=\mathcal{D}_{cc}.$
The proof of Lemma 8 is given in Appendix I.
Theorem 4.
Under Assumption 1, 2, and 3, and considering the case that $P_{k-1|k-1}=P_{cc,k-1|k-1}$, one has
$\lim_{l\to+\infty}P_{i,k|k}=P_{c,k|k}.$ (47)
The proof of Theorem 4 is given in Appendix J.
Remark 11.
Based on the structural characteristics of $\mathcal{A}(l)$, $\mathcal{B}(l)$, and $\mathcal{D}(l)$, Lemma 8 shows that the matrices $\mathcal{A}(l)$, $\mathcal{B}(l)$, and $\mathcal{D}(l)$ approach to the centralized matrices $\mathcal{A}_{cc}$, $\mathcal{B}_{cc}$, and $\mathcal{D}_{cc}$, respectively, as the fusion step tends to infinity. Then, it is demonstrated that the transient performance exhibits similar results. In other words, as the fusion step increases, the performance of the proposed filter approaches that of the centralized filter. By considering both the transient properties and the convergence properties, we can gain comprehensive insights into the performance of the proposed algorithm.
5 Simulations
In this section, the effectiveness of the theoretical results is validated through a target tracking numerical experiment. Consider a sensor network comprising five sensors labeled from 1 to 5, and its communication topology is illustrated in Fig. 1.
Figure 1: The diagram of the communication topology.
Consider the target dynamics described by
$A=\left(\begin{array}[]{cc}I_{2}&TI_{2}\\ 0&I_{2}\end{array}\right),$ (48)
where $T=0.25$ is the discretization interval. The process noise covariance is defined as
$\bar{Q}=\left(\begin{array}[]{cc}\frac{T^{3}}{3}&\frac{T^{2}}{2}\\ \frac{T^{2}}{2}&T\end{array}\right),~{}~{}~{}Q=\left(\begin{array}[]{cc}\bar{Q}&0.5\bar{Q}\\ 0.5\bar{Q}&\bar{Q}\end{array}\right).$ (49)
Two kinds of sensors are employed in the sensor network: position sensors and velocity sensors. The observation matrix for the position sensors is given by
$C_{p}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\end{array}\right),$ (50)
with the measurement noise covariance $R_{p}=\text{diag}\{1,1\}$. Similarly, the observation matrix for the velocity sensors is represented by
$C_{v}=\left(\begin{array}[]{cccc}0&0&1&0\\ 0&0&0&1\end{array}\right),$ (51)
with the measurement noise covariance $R_{v}=\text{diag}\{5,5\}$. In this sensor network, it is assumed that sensor 1, 2, and 4 are the position sensors, while sensor 3 and 5 serve as the velocity sensors. The initial state is set as $x_{0}=[1;1;1;1]$, and the initial state estimate is set as a random variable with the mean $\hat{x}_{i,k|k}=x_{0}$ and the initial estimation covariance $P_{i,0|0}=\text{diag}\{10,10,10,10\}$. The parameter $\mu_{ij}$ is selected as $\frac{1}{l_{ii}+a_{il}+1}$.
The mean square error (MSE) is utilized to evaluate the performance of the estimator based on the Monte Carlo method, described by
$MSE_{i,k}=\frac{1}{M}\sum^{M}_{l=1}\|\hat{x}^{(l)}_{i,k}-x^{(l)}_{k}\|^{2}_{2},$ (52)
where $M$ is the trial number, and $\hat{x}^{(l)}_{i,k}$ and $x^{(l)}_{k}$ are the state estimate and the true state at the $l$-th trial, respectively. In the simulation, $M$ is set as $1000$. All simulations are executed in MATLAB R2020a with an Intel Core i5-1135G7 CPU @ 2.40 GHz.
Two examples are designed to verify the effectiveness of the proposed algorithm. Example 1 aims to demonstrate the steady-state performance of the proposed COMDF as the fusion step $l$ increases. Example 2 is designed to show the performances and properties of COMDF in comparison with other existing consensus-based distributed filers.
Figure 2: Illustration figure for the steady-state performance of COMDF with the increasing fusion step $l$.
Example 1: Fig. 2 exhibits the steady-state performance of COMDF of five sensors with the increasing fusion step $l$. It is shown that the performance gap between the centralized filter and COMDF is exponential convergence (Theorem 3). In addition, a small fusion step can also ensure the performance of COMDF.
Example 2: To assess the performance of the proposed COMDF, three other algorithms are considered: the consensus-on-measurement distributed filter (CMDF) from [6], the consensus-on-information distributed filter (CIDF) from [4], and the centralized Kalman filter (CKF) in Section 4.1. For comparison, it is assumed that the communication topology is undirected in Fig. 1, and the fusion step is set as $l=10$.
Fig. 3 displays the performance of four algorithms with the increasing time step $k$. As the time step $k$ tends to infinity, four algorithms converge. CKF and COMDF are the algorithms that fix the gain matrix $K$, hence, they may not be optimal in every step. However, as $k$ tends to infinity, CKF approaches optimality. Meanwhile, the performance of COMDF approaches that of CKF with the increasing fusion step. Consequently, before the algorithms converge, CIDF and CMDF show a faster convergence velocity. CIDF adopts the covariance intersection method to handle the noise correlations, and the steady-state performance is degraded.
Figure 3: Illustration figure for the performance of four algorithms with the increasing time step $k$.
Table 1 presents the time consumption and the memory usage of four algorithms in Example 2. All four algorithms exhibit similar memory usage, but substantial differences emerge in terms of time consumption. Notably, COMDF demonstrates the smallest time consumption among the three distributed algorithms, attributed to its transmission of only measurements. Conversely, CIDF exhibits the highest time consumption among the three distributed filters, possibly due to a higher frequency of the inverse operation in the algorithm.
Table 1: The time consumption and the memory usage of four algorithms in Example 2.
Algorithm Time Consumption (s) Memory Usage (MB)
COMDF      23.864      2192
CMDF      32.709      2183
CIDF      144.438      2189
CKF      2.005      2194
6 Conclusions
This paper proposes a consensus-on-measurement distributed filter over directed graphs, embedded with an augmented leader-following measurement fusion strategy. The parameters are designed to guarantee the uniformly upper bound of the estimation error covariances. The steady-state and transient performances are analyzed with the increasing fusion step, and the relations between the proposed distributed filter and the centralized filter are revealed. In the future, it is desired to design a dynamic gain matrix to optimize the distributed filter at every step, aiming for a faster convergence. Additionally, there is an intention to explore a new parameter design method to reduce the spectral radius of the communication matrix.
Appendix A PROOF of Proposition 1
Item 1): Using equation (11), $\varepsilon^{(l)}_{i,k}$ can be derived as
$\varepsilon^{(l)}_{i,k}$ $=z^{(l-1)}_{i,k}-\Lambda_{i}\Big{[}\sum^{N}_{l=1}a_{il}(z^{(l-1)}_{i,k}-z^{(l-1)}_{l,k})$ (53)
$~{}~{}~{}~{}+B_{i}(z^{(l-1)}_{i,k}-y_{k})\Big{]}-y_{k}$
$=\varepsilon^{(l-1)}_{i,k}-\Lambda_{i}\Big{[}\sum^{N}_{l=1}a_{il}(\varepsilon^{(l-1)}_{i,k}-\varepsilon^{(l-1)}_{l,k})+B_{i}\varepsilon^{(l-1)}_{i,k}\Big{]}.$
Then, the augmented vector $\varepsilon^{(l)}_{k}$ can be obtained as
$\varepsilon^{(l)}_{k}$ $=G\varepsilon^{(l-1)}_{k},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (54)
where $G$ is given in (13). When $l=0$, it follows
$\varepsilon^{(0)}_{i,k}$ $=z^{(0)}_{i,k}-y_{k}$ (55)
$=CAe_{i,k-1}-C\omega_{k-1}-\nu_{k}.$
Similarly, the augmented vector $\varepsilon^{(0)}_{k}$ is
$\varepsilon^{(0)}_{k}$ $=(I_{N}\otimes CA)e_{k-1}-(I_{N}\otimes C)(1_{N}\otimes\omega_{k-1})$ (56)
$~{}~{}~{}~{}-1_{N}\otimes\nu_{k}.$
Next, it can be concluded that
$\varepsilon^{(l)}_{k}$ $=G\varepsilon^{(l-1)}_{k}$ (57)
$=G^{l}\varepsilon^{(0)}_{k}.$
Item 2): Based on the results of Item 1), $E\{\varepsilon^{(l)}_{k}\}$ and $E\{\varepsilon^{(l)}_{k}(\varepsilon^{(l)}_{k})^{T}\}$ can be calculated.
Appendix B PROOF of Lemma 2
According to Lemma 2, since $\rho(G)<1$, it follows $\lim_{l\to+\infty}G^{l}=0$. Observing (17) and (18), $G^{l}$ exists in all terms. Hence, this lemma can be proven.
Appendix C PROOF of Proposition 2
Utilizing the state estimator equations (4) and (5) in conjunction with the dynamical system (1), the estimation error $e_{i,k}$ for sensor $i$ is
$e_{i,k}$ $=(A-KCA)\hat{x}_{i,k-1|k-1}+Kz^{(l)}_{i,k}-Ax_{k-1}-\omega_{k-1}$
$=(A-KCA)e_{i,k-1}+K\varepsilon^{(l)}_{i,k}$
$~{}~{}~{}~{}-(I-KC)\omega_{k-1}+K\nu_{k}.$
Subsequently, by using Proposition 1, the augmented estimation error $e_{k}$, i.e., $e_{k}=[e^{T}_{1,k},\ldots,e^{T}_{N,k}]^{T}$, can be obtained as
$e_{k}$ $=(I_{N}\otimes(A-KCA))e_{k-1}$ (58)
$~{}~{}~{}~{}+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA)e_{k-1}$
$~{}~{}~{}~{}-(I_{N}\otimes K)G^{l}(1_{N}\otimes(C\omega_{k-1}+\nu_{k}))$
$~{}~{}~{}~{}+1_{N}\otimes(-(I-KC)\omega_{k-1}+K\nu_{k}).$
Note that $1_{N}\otimes(K\nu_{k})=(I_{N}\otimes K)(1_{N}\otimes\nu_{k})$ and $1_{N}\otimes(C\omega_{k-1})=(I_{N}\otimes C)(1_{N}\otimes\omega_{k-1})$, and (58) can be rewritten as
$e_{k}$ $=(I_{N}\otimes(A-KCA)+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA))e_{k-1}$
$~{}~{}~{}~{}-(I_{N}\otimes(I-KC)-(I_{N}\otimes K)$
×$~{}~{}~{}~{}\times G^{l}(I_{N}\otimes C))(1_{N}\otimes\omega_{k-1})$
$~{}~{}~{}~{}+(I_{N}\otimes K)(I_{Nr}-G^{l})(1_{N}\otimes\nu_{k}).$
By denoting the matrices as (24), (25), and (26), $e_{k}$ has the following form
$e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}).$
Finally, $P_{k|k}$ can be calculated as (22) and (23) according to $P_{k|k}=E\{e_{k}e^{T}_{k}\}$
Appendix D PROOF of Lemma 6
Under the condition $0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}}$ and the fact $l_{ij}\leq 0,i\neq j$, it holds
$1-\mu_{il}(l_{ii}+a_{il})\geq 0$ (59)
and
$-\mu_{il}l_{ij}\geq 0.$ (60)
By combining (27), (28), (59), and (60), it can be concluded that the matrix $G$ is nonnegative. By utilizing Lemma 3 and observing (29), it follows
$\rho(G)$ $\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}G_{ij}$
$\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}g_{[ij,l]}$
$\leq 1-\mu_{il}a_{il}.$
Hence, the conclusion is drawn that $\rho(G)\leq 1$.
Subsequently, it will be demonstrated that the eigenvalues of $G$ do not include the value $1$. By substituting the value $1$ into the characteristic polynomial of $G$, one has
$|I-G|=|\Lambda(\mathcal{L}\otimes I_{r}+B)|.$
According to Definition 1, $L_{B}=\mathcal{L}\otimes I_{r}+B$ is irreducibly diagonally dominant. Then, by utilizing Lemma 1, $L_{B}=\mathcal{L}\otimes I_{r}+B$ is nonsingular. Additionally, $\Lambda$ is positive definite with $\text{rank}(\Lambda)=Nr$, since $\mu_{ij}>0$. Utilizing Sylvester’s rank inequality, one obtains $\text{rank}(\Lambda)+\text{rank}(L_{B})-Nr\leq\text{rank}(\Lambda L_{B})$ and $\text{rank}(\Lambda L_{B})=Nr$. Hence, 1 is not an eigenvalue of the matrix $G$. In conclusion, it can be inferred that $\rho(G)<1$.
Appendix E PROOF of Theorem 1
Since $\|G\|_{2}<1$, one can conclude that $G^{l}$ is uniformly upper-bounded for any positive $l$. Considering (25) and (26), both $\mathcal{B}(l)$ and $\mathcal{D}(l)$ are also uniformly upper-bounded due to the boundedness of $K$ and $C$. Likewise, it can be deduced that $\Phi(l)$ in (23) is uniformly upper-bounded by using the boundedness of $Q$ and $R$. If $\mathcal{A}(l)$ is Schur stable, it can be concluded that $P_{k|k}$ is uniformly upper-bounded.
Given Assumption 2, it is well known that the matrix $A-KCA$ is Schur stable [2]. Applying Lemma 2, as $l$ tends to infinity, $\lim_{l\to\infty}G^{l}\to 0$. Consequently, one observes $\lim_{l\to\infty}(I_{N}\otimes K)G^{l}(I_{N}\otimes CA)\to 0$. It is no doubt that there exists a positive $l$ such that $\mathcal{A}(l)$ is Schur stable.
In matrix theory, it is established that the spectral radius of $\rho(\mathcal{A}(l))$ is bounded by the spectral norm $\|\mathcal{A}(l)\|_{2}$, i.e., $\rho(\mathcal{A}(l))\leq\|\mathcal{A}(l)\|_{2}$. Using the identity $\|J_{1}\otimes J_{2}\|_{2}=\|J_{1}\|_{2}\|J_{2}\|_{2}$, it follows
$\|\mathcal{A}(l)\|_{2}\leq\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}.$ (61)
Let $\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}<1$. After some algebraic manipulations, one has
$\|G\|^{l}_{2}<\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}.$ (62)
Since $10<\|G\|_{2}<$ 1, there exists
$l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}},$ (63)
such that when $l>l_{0}$, $\mathcal{A}(l)$ is Schur stable. Consequently, it can be proven that $P_{k|k}$ is uniformly upper-bounded. Since
$P_{i,k|k}$ $=[0_{n\times n(i-1)},I_{n},0_{n\times n(N-i)}]P_{k|k}$ (64)
×$~{}~{}~{}~{}\times[0^{T}_{n\times n(i-1)},I^{T}_{n},0^{T}_{n\times n(N-i)}]^{T},$
it can be found that $P_{i,k|k}$ is also uniformly upper-bounded.
Appendix F PROOF of Theorem 2
Under Assumption 1, 2, and 3, it can be concluded that $\mathcal{A}(l)$ and $\mathcal{A}_{cc}$ are Schur stable. In addition, $\Phi(l)$ and $\Phi_{cc}$ are also uniformly upper-bounded due to the boundedness of $K$, $C$, $Q$, $R$, and $G^{l}$. By utilizing Theorem 1 in [7], it can be proven that $P_{k|k}$ and $P_{cc,k|k}$ converge to $P^{(l)}$ and $P_{cc}$, respectively.
Appendix G PROOF of Lemma 7
First, consider the term $\mathcal{A}(l)P^{(l)}\mathcal{A}^{T}(l)$ in $P^{(l)}$, and one has
$\mathcal{A}(l)P^{(l)}\mathcal{A}^{T}(l)$ $=(\mathcal{A}_{cc}+{\mathcal{\bar{A}}}(l))P^{(l)}(\mathcal{A}_{cc}+\bar{\mathcal{A}}(l))^{T}$ (65)
$=\mathcal{A}_{cc}P^{(l)}\mathcal{A}^{T}_{cc}+{\mathcal{\bar{A}}}(l)P^{(l)}\mathcal{A}^{T}_{cc}$
$~{}~{}~{}+\mathcal{A}_{cc}P^{(l)}{\mathcal{\bar{A}}}^{T}(l)+{\mathcal{\bar{A}}}(l)P^{(l)}{\mathcal{\bar{A}}}^{T}(l).$
By utilizing (65), $P^{(l)}-P_{cc}$ has the following form
$P^{(l)}-P_{cc}=\mathcal{A}_{cc}(P^{(l)}-P_{cc})\mathcal{A}^{T}_{cc}+\bar{\Phi}(l),$ (66)
where
$\bar{\Phi}(l)$ $={\mathcal{\bar{A}}}(l)P^{(l)}\mathcal{A}^{T}_{cc}+\mathcal{A}_{cc}P^{(l)}{\mathcal{\bar{A}}}^{T}(l)+{\mathcal{\bar{A}}}(l)P^{(l)}{\mathcal{\bar{A}}}^{T}(l)$ (67)
$~{}~{}~{}+\Phi(l)-\Phi_{cc}.$
Performing the iteration (66) for $k$ times, it follows
$P^{(l)}-P_{cc}$ $=\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}$ (68)
$~{}~{}~{}~{}+\sum^{k-1}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$
By performing an infinite number of iterations and utilizing the fact that $\mathcal{A}_{cc}$ is Schur stable, it holds
$\lim_{k\to\infty}\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}=0$
and
$P^{(l)}-P_{cc}=\sum^{\infty}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$
Next, similarly to (65), $\bar{\Phi}(l)$ in (67) can be calculated as (44).
Appendix H PROOF of Theorem 3
By calculating the spectral norm of (43), one has
$\|P^{(l)}-P_{cc}\|_{2}\leq\|\bar{\Phi}(l)\|_{2}\sum^{\infty}_{k=0}\|\mathcal{A}^{k}_{cc}\|^{2}_{2}.$ (69)
For $\|\bar{\Phi}(l)\|_{2}$, based on (44), it follows
$\|\bar{\Phi}(l)\|_{2}$ $\leq(\|{\mathcal{\bar{A}}}(l)\|_{2}+2\|\mathcal{A}_{cc}\|_{2})\|P^{(l)}\|_{2}\|{\mathcal{\bar{A}}}(l)\|_{2}$ (70)
[…]
$~{}~{}~{}+N(\|{\mathcal{\bar{D}}}(l)\|_{2}+2\|\mathcal{D}_{cc}\|_{2})\|R\|_{2}\|{\mathcal{\bar{D}}}(l)\|_{2}.$
[…]
$\|\bar{\Phi}(l)\|_{2}\leq M_{3}\|G^{l}\|_{2}$ (71)
holds, where
$M_{3}$ $=(\|{\mathcal{\bar{A}}}(l)\|_{2}+2\|\mathcal{A}_{cc}\|_{2})\|P^{(l)}\|_{2}\|K\|_{2}\|CA\|_{2}$ (72)
[…]
$~{}~{}~{}+N(\|{\mathcal{\bar{D}}}(l)\|_{2}+2\|\mathcal{D}_{cc}\|_{2})\|R\|_{2}\|K\|_{2}.$
By using Lemma 4 and the similar technique in [26], a positive number $M_{4}$ can be found such that
$\|\mathcal{A}^{k}_{cc}\|_{2}$ $\leq\sqrt{n}\sum^{n-1}_{j=0}\binom{n-1}{j}\binom{k}{j}\|\mathcal{A}_{cc}\|^{j}_{2}\rho(\mathcal{A}_{cc})^{k-j}$ (73)
$\leq\sqrt{n}\Bigg{(}n\text{max}_{j}\binom{n-1}{j}\|\mathcal{A}_{cc}\|^{j}_{2}\Bigg{)}k^{n}\rho(\mathcal{A}_{cc})^{k-n}$
$\leq M_{4}k^{n}\rho(\mathcal{A}_{cc})^{k-n}.$
Since $\rho(\mathcal{A}_{cc})<1$, the convergence of the infinite sum $\sum^{\infty}_{k=0}\|\mathcal{A}^{k}_{cc}\|_{2}$ can be proven. Since $\|\mathcal{A}^{k}_{cc}\|_{2}$ is uniformly bounded for all $k$, there exists a positive number $M_{5}$ such that
$\sum^{\infty}_{k=0}\|\mathcal{A}^{k}_{cc}\|^{2}_{2}\leq M_{5}.$ (74)
Next, the term $\|G^{l}\|_{2}$ is analyzed, and the results of two items are proven respectively.
Item 1): Similarly to (73), there exists a positive number $M_{6}$ such that
$\|G^{l}\|_{2}$ $\leq\sqrt{Nr}\sum^{Nr-1}_{j=0}\binom{Nr-1}{j}\binom{l}{j}\|G\|^{j}_{2}\rho(G)^{l-j}$ (75)
$\leq M_{6}l^{Nr}\rho(G)^{l-Nr}.$
By combining (69), (71), (74), and (75), a positive number $M_{1}$ can be designed such that
$\|P^{(l)}-P_{cc}\|_{2}\leq M_{1}l^{Nr}\rho(G)^{l-Nr}.$ (76)
Item 2): If $\|G\|_{2}<1$, (71) can be calculated as
$\|\bar{\Phi}(l)\|_{2}\leq M_{3}\|G\|^{l}_{2}.$ (77)
By combining (69), (74), and (77), a positive number $M_{2}$ can be obtained such that
$\|P^{(l)}-P_{cc}\|_{2}\leq M_{2}\|G\|^{l}_{2}.$ (78)
Appendix I PROOF of Lemma 8
Based on Assumption 3 and Lemma 2, as $l$ tends to infinity, one has $\lim_{k\to\infty}G^{l}=0$. By combining (24), (25), and (26) and comparing them to (38), (39), and (40), this lemma can be proven.
Appendix J PROOF of Theorem 4
Utilizing Lemma 8, (22), and (36), it follows
$\lim_{l\to+\infty}P_{k|k}=P_{cc,k|k}.$ (79)
Based on (64), this conclusion can be drawn.

### v2

@@
-Designing Consensus-Based Distributed Filtering over Directed Graphs
+Distributed State Estimation for Discrete-Time Linear Systems over Directed Graphs: A Measurement Perspective
 Abstract
-This paper proposes a novel consensus-on-only-measurement distributed filter over directed graphs under the collectively observability condition.
-First, the distributed filter structure is designed with an augmented leader-following measurement fusion strategy.
-Subsequently, two parameter design methods are presented, and the consensus gain parameter is devised utilizing local information exclusively rather than global information.
-Additionally, the lower bound of the fusion step is derived to guarantee a uniformly upper bound of the estimation error covariance.
-Moreover, the lower bounds of the convergence rates of the steady-state performance gap between the proposed algorithm and the centralized filter are provided with the fusion step approaching infinity.
-The analysis demonstrates that the convergence rate is, at a minimum, as rapid as exponential convergence under the spectral norm condition of the communication graph.
-The transient performance is also analyzed with the fusion step tending to infinity.
-The inherent trade-off between the communication cost and the filtering performance is revealed from the analysis of the steady-state performance and the transient performance.
-Finally, the theoretical results are substantiated through the validation of two simulation examples.
-keywords:
-Distributed filtering; Consensus; Sensor networks; Performance analysis; Directed graph.
-1 Introduction
+This paper proposes a novel consensus-based distributed filter over directed graphs under the collectively observability condition.
+The distributed filter is designed using an augmented leader-following information fusion strategy, and the gain parameter is determined exclusively using local information.
+Additionally, the lower bound of the fusion step number is derived to ensure that the estimation error covariance remains uniformly upper-bounded.
+Furthermore, the lower bounds for the convergence rates of the steady-state performance gap between the proposed filter and the centralized filter are provided as the fusion step number approaches infinity.
+The analysis demonstrates that the convergence rate is at least as fast as exponential convergence, provided the communication topology satisfies the spectral norm condition.
+Finally, the theoretical results are validated through two simulation examples.
+I Introduction
 In the preceding two decades, investigations into sensor networks have occupied a prominent position in the realm of systems and control, owing to its extensive applications encompassing health care monitoring [13, 1], environmental sensing [7, 10], and collaborative mapping [9, 18].
 In order to meet the demands for the network reliability and alleviate the computational and communicative burdens on energy-constrained sensors, the implementation of distributed state estimation serves as a good solution to these challenges.
-The objective of distributed state estimation is to estimate the state of the target system for each sensor in sensor networks by utilizing information from local neighbors, including local measurements, local state estimates, or related data.
-Founded on various dynamical systems and constraint conditions, a wealth of research achievements regarding the distributed state estimation have emerged.
-The evolution of consensus theory within the domain of multi-agent systems has imparted profound insights into the information fusion for distributed state estimation [24, 15, 29, 20].
-Several reputable consensus-based approaches to distributed state estimation have been advanced [22, 23, 7, 4, 6, 30].
-In [21, 22], consensus filters, including low-pass, high-pass, and band-pass filters, were employed to fuse the covariances, measurements, or state estimates for both continuous-time and discrete-time systems.
-In [23], consensus terms were designed in the filter structure, and the optimal gain was derived following with a Lyapunov-based stability analysis.
-Diffusion strategies, where each sensor transmitted the information to its neighbors, were proposed for distributed Kalman filtering and smoothing in [7].
-In [4], consensus of the normalized geometric mean of the probability density functions, called the Kullback-Leibler average, was proposed, and this fusion rule aligned with the concepts of the generalized covariance intersection.
-In [6], the consensus algorithms on measurements and information were combined to preserve the positive attributes of both approaches, thereby enhancing energy efficiency while diminishing the conservative assumption regarding noise correlation.
-A distributed state estimation filter by using consensus-based information fusion strategies for continuous-time systems with correlated measurement noise was proposed in [10].
-The applicability of distributed algorithms is influenced by the constraints imposed on the communication topologies.
-There usually exist two key assumptions: undirected connected graphs [22, 9, 28, 10, 19] and directed and strongly connected graphs [4, 6, 30].
-In general, directed graphs can model more intricate communication topologies compared to undirected graphs.
-For directed graphs, the works [4, 6] stated that the consensus matrix needed to be primitive and doubly stochastic to achieve an average consensus, with the introduction of Metropolis weights under the assumption that the graph is undirected.
-However, challenges arise when applying these principles to directed graphs.
-Hence, this paper focuses on designing algorithms and their corresponding parameters over directed graphs characterized by an adjacent matrix with elements 00 and $1$.
-The classical distributed Kalman filter structure is characterized by the iterations of vectors and covariance matrices.
-To avoid the computation and transmission of numerous matrices, a consensus-on-only-measurement strategy is considered in this paper.
-Furthermore, the detailed performance of the proposed distributed filter is analyzed.
-There exist various performance analysis methods for distributed filters.
-The Lyapunov’s second method for stability was utilized in many works [23, 5, 4, 6] by constructing the Lyapunov function and conducting stability analysis.
-Recently, the performance of the distributed filter was analyzed by investigating the difference between the solutions of a modified discrete-time algebraic Riccati equation and a discrete-time Lyapunov equation in [26].
-For distributed filters, the performance gap between the distributed filter and the centralized filter is a significant concern, and this paper also focuses on this aspect.
-Motivated by the aforementioned observations, this paper aims to propose a consensus-on-only-measurement distributed filter over directed graphs, and furnish comprehensive performance analysis of the proposed distributed filter.
+The objective of distributed state estimation is to estimate the state of the target system for each sensor in sensor networks.
+Building on various dynamical systems and constraint conditions, a wealth of research has emerged in the field of distributed state estimation.
+The evolution of consensus theory within the domain of multi-agent systems has imparted profound insights into the information fusion for distributed state estimation [15, 12, 22].
+Several reputable consensus-based approaches to distributed state estimation have been proposed [16, 23, 3, 5, 17, 4, 8].
+Depending on the filter structures, these filters fuse different types of information from local neighbors, including local state estimates [5, 23, 3], intermediate quantities [17, 4], or measurements [8].
+For the fusion of local state estimates, Battistelli et al.
+[5] introduced a distributed filter based on the covariance intersection method, which combines estimates from different sensors.
+In another approach, Battilotti et al.
+[3] developed a distributed filter that fuses the state estimation deviations among neighboring sensors.
+Regarding the fusion of intermediate quantities, Olfati-Saber [17] designed a Kalman-consensus filter using intermediate quantities to approximate the centralized filter, with a Lyapunov-based stability analysis.
+Additionally, Battistelli et al.
+[4] presented a distributed filter that fuses both state estimates and intermediate quantities, thereby preserving the benefits of both approaches.
+For the fusion of local measurements, Duan et al.
+[8] proposed a distributed state estimation filter for continuous-time systems with correlated measurement noise.
+However, compared to existing literature, the research on the fusion of local measurements has not been thoroughly investigated for discrete-time linear systems due to the following challenges:
+  1. 1.
+    Designing the filter structure and information fusion strategies using local measurements is challenging while ensuring the filter performance.
+  2. 2.
+    Existing gain design methods for directed graphs [5, 4] face complexities in applications, and developing an effective distributed design method is crucial.
+  3. 3.
+    It is necessary to explore the performance gap between the proposed filter and the centralized filter, with a particular focus on the convergence rate regarding the fusion step number.
+Motivated by the above observations, this paper aims to propose a novel consensus-based distributed filter over directed graphs using a dynamic averaging algorithm, and provides a comprehensive performance analysis of the proposed distributed filter.
 The main contributions of this paper are summarized below:
   1. 1.
-    A novel consensus-on-only-measurement distributed filter (COMDF) over directed graphs is proposed for discrete-time systems under the collectively observability condition, and an augmented leader-following measurement fusion strategy is presented to estimate the other sensors’ measurements.
-Transmitting only measurements eliminates the need for the computation and transmission of numerous matrices.
-Additionally, the properties of the measurement estimate error and the state estimate error are presented (Proposition 1 and 2).
+    We propose a novel consensus-based distributed filter using an augmented leader-following information fusion strategy over directed graphs under the collectively observability condition.
+Compared to the existing literature [16, 23, 3, 5, 17, 4, 8], this filter provides a novel filter structure that has been rarely studied.
   2. 2.
-    The parameter design methods are provided, including a distributed design method for each sensor’s consensus gain and the establishment of a lower bound of the fusion step (Theorem 1).
-Compared to the unified design method [10], the proposed distributed design method exclusively relies on local information, facilitating the distributed usage without the need for global information.
-The lower bound of the fusion step illuminates the necessary conditions to guarantee the uniformly upper bound of the estimation error covariance.
+    Under the coupling between the measurement estimation error and the state estimation error, we provide gain design methods that rely solely on local information.
+Additionally, we derive the lower bound for the fusion step number (Theorem 1).
   3. 3.
-    The impact of the fusion step on the steady-state performance of COMDF is analyzed under the conditions involving the spectral radius and the spectral norm of the communication matrix.
-The lower bounds of the convergence rates of the steady-state performance gap between COMDF and the centralized filter are provided with the fusion step approaching infinity.
-Particularly, under the spectral norm condition, the convergence rate is at least as fast as exponential convergence (Theorem 3).
-Additionally, the transient performance is also discussed with the fusion step tending to infinity (Theorem 4).
+    We analyze the impact of the fusion step number on the steady-state performance.
+Specifically, we provide the lower bounds on the convergence rates of the steady-state performance gap between the proposed distributed filter and the centralized filter as the fusion step approaches infinity (Theorem 3).
 The remainder of this paper is organized as follows.
-In Section 2, some preliminaries, including graph concepts, system models, and some useful lemmas, and the problem formulation are provided.
-In Section 3, a consensus-based distributed filter is proposed over directed graphs, and two parameter design methods are given.
-In Section 4, the steady-state and the transient performances of the proposed filter are analyzed with the increasing fusion step.
-In Section 5, two numerical examples are provided to validate the effectiveness of the obtained results.
-In Section 6, conclusions are drawn.
-Notations: Throughout this paper, define $\mathcal{R}^{n}$ as the sets of $n$-dimensional real vectors and $\mathcal{R}^{n\times n}$ as $n\times n$-dimensional real matrices.
-For a matrix $A\in\mathcal{R}^{n\times n}$, let $A^{-1}$ and $A^{T}$ represent its inverse and transpose, respectively, $\|A\|_{2}$ is the spectral norm, $\rho(A)$ is the spectral radius, and $[A]_{ij}$ denotes the $(i,j)$-th element of the matrix $A$.
-Notation $\otimes$ represents the Kronecker product.
-The matrix inequalities $A>B$ and $A\geq B$ signify that $A-B$ is positive definite and positive semi-definite, respectively.
-$E\{x\}$ denotes the expectation of the random variable $x$.
-2 Preliminaries and Problem Statement
-2.1 Graph Theory
-The communication topology ${\mathcal{G}}(\mathcal{V},\mathcal{E})$ is utilized to illustrate the nodes and the communication links within a sensor network, where the node set $\mathcal{V}=\{1,2,\ldots,N\}$ and the edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$.
-For $i,j\in\mathcal{V}$, $(j,i)$ signifies that node $j$ can transmit information to node $i$, and node $j$ is called a neighbor of node $i$.
-The neighbor set of node $i$ is denoted as $\mathcal{N}_{i}=\{j|(j,i)\in\mathcal{V}\}$, and $|\mathcal{N}_{i}|$ is the cardinality of the neighbors of node $i$.
-The adjacent matrix is $S=[a_{ij}]_{N\times N}$, where $a_{ij}=1$ if $(j,i)\in\mathcal{E}$ and $a_{ij}=0$ otherwise.
-Let $D=\text{diag}\{|\mathcal{N}_{1}|,\ldots,|\mathcal{N}_{N}|\}$, and the Laplacian matrix is defined as $\mathcal{L}=D-S$.
-The edge $(i,j)$ is undirected if $(i,j)\in\mathcal{E}$ implies $(j,i)\in\mathcal{E}$.
-The communication graph is termed undirected if every edge is undirected.
-The graph $G$ contains a directed path from node $i_{1}$ to node $i_{m}$, if there exists a sequence of connected edges $(i_{k},i_{k+1}),k=1,\ldots,m-1$.
-The communication graph is called strongly connected, if there exists a path between any pair of distinct nodes.
-Assumption 1.
-The communication graph is directed and strongly connected.
-Remark 1.
-Assumption 1 can be extended to a jointly connected switching topology by applying the corresponding definition of communication networks [27, 17, 18].
-The proposed parameter design methods are applicable in these situations as well.
-For simplicity, this paper focuses on the directed and strongly connected topology assumption.
-2.2 System Model
+Section II presents the preliminaries and problem formulation.
+Section III proposes a consensus-based distributed filter over directed graphs, and provides two parameter design methods.
+Section IV analyzes the steady-state performance of the proposed filter as the fusion step number increases.
+Section V presents two numerical examples to validate the effectiveness of the results.
+Finally, Section VI concludes the paper.
+Notations: For a matrix $A\in\mathcal{R}^{n\times n}$, $\|A\|_{2}$ is the spectral norm, $\rho(A)$ is the spectral radius, and $A]_{ij}$ denotes the $(i,j)$-th element of the matrix $A$.
+Symbol $\otimes$ represents the Kronecker product.
+The matrix inequalities $A>B$ and $A\geq B$ signify that ABA-B is positive definite and positive semi-definite, respectively.
+II Preliminaries and Problem Statement
+II-A Some Useful Lemmas
+Definition 1
+[11] For a matrix $\in\mathcal{R}^{n\times n}, MM$ is said to be irreducibly diagonally dominant if 1) $M$ is irreducible, 2) $M$ is diagonally dominant, i.e., $|m_{ii}|\geq R^{s}_{i}(M)$ for all $i=1,\ldots,n$, 3) There is an $i\in\{1,\ldots,n\}$ such that $|m_{ii}|>R^{s}_{i}(M)$, where $R^{s}_{i}(M)=\sum_{j\neq i}|m_{ij}|$.
+Lemma 1
+[11] Let the matrix $M$ be irreducibly diagonally dominant. Then, 1) $M$ is nonsingular, 2) If $M$ is Hermitian and every main diagonal entry is positive, $M$ is positive definite.
+Lemma 2
+[11] For a matrix $M$ and a positive integer $k$, it holds $\lim_{k\to\infty}M^{k}=0$ if and only if $\rho(M)<1$.
+Lemma 3
+[11] For a nonnegative matrix $M=[m_{ij}]$, it holds $\text{min}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}\leq\rho(M)\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}.$
+Lemma 4
+[19] For any matrix $M\in\mathcal{R}^{n\times n}$, it holds $\|M^{k}\|_{2}\leq\sqrt{n}\sum^{n-1}_{j=0}\binom{n-1}{j}\binom{k}{j}\|M\|^{j}_{2}\rho(M)^{k-j},$ where $\binom{m}{n}$ is the combinatorial number with $\binom{k}{j}=0$ for $j>k$.
+II-B System Model
 Consider a discrete-time linear time-invariant system observed by a network of $N$ sensors:
 $x_{k+1}=Ax_{k}+\omega_{k},$ (1)
-$y_{i,k}=C_{i}x_{k}+\nu_{i,k},~{}~{}~{}i=1,2,...,N,$
-where $k$ is the discrete-time index, $i\in\mathcal{V}$ is the $i$-th sensor of the network, $x_{k}\in\mathcal{R}^{n}$ is the system state vector, $y_{i,k}\in\mathcal{R}^{r_{i}}$ is the measurement vector taken by sensor $i$, $A\in\mathcal{R}^{n\times n}$ is the state transition matrix, $C_{i}\in\mathcal{R}^{r_{i}\times n}$ is the observation matrix of sensor $i$, and $\omega_{k}\in\mathcal{R}^{n}$ and $\nu_{i,k}\in\mathcal{R}^{r_{i}}$ are the zero-mean Gaussian noise with the covariances $Q_{k}\in\mathcal{R}^{n\times n}$ and $R_{i,k}\in\mathcal{R}^{r_{i}\times r_{i}}$, respectively.
+$y_{i,k}=C_{i}x_{k}+\nu_{i,k},~~~i=1,2,.,N,$
+where $k$ represents the discrete-time index, $i$ represents the $i$-th sensor in the network, $x_{k}\in\mathcal{R}^{n}$ is the system state vector, $y_{i,k}\in\mathcal{R}^{r_{i}}$ is the measurement vector of sensor $i$, $A\in\mathcal{R}^{n\times n}$ is the state transition matrix, $C_{i}\in\mathcal{R}^{r_{i}\times n}$ is the observation matrix, and $\omega_{k}\in\mathcal{R}^{n}$ and $\nu_{i,k}\in\mathcal{R}^{r_{i}}$ are zero-mean Gaussian noise with the covariances $Q_{k}\in\mathcal{R}^{n\times n}$ and $R_{i,k}\in\mathcal{R}^{r_{i}\times r_{i}}$, respectively.
 The noise sequences $\{\omega_{k}\}^{\infty}_{k=0}$ and $\{\nu_{i,k}\}^{\infty,N}_{k=0,i=1}$ are mutually uncorrelated.
-For the whole network, $C=[C^{T}_{1},\ldots,C^{T}_{N}]^{T}\in\mathcal{R}^{r\times n}$ is the augmented observation matrix, and $R=\text{diag}\{R_{1},\ldots,R_{N}\}\in\mathcal{R}^{r\times r}$ is the augmented measurement noise covariance matrix, where $r=\sum^{N}_{i=1}r_{i}$.
-Assumption 2.
+Let $C=[C^{T}_{1},\ldots,C^{T}_{N}]^{T}\in\mathcal{R}^{r\times n}$ denote the augmented observation matrix, and $R=\text{diag}\{R_{1},\ldots,R_{N}\}\in\mathcal{R}^{r\times r}$ represent the augmented measurement noise covariance matrix, where $r=\sum^{N}_{i=1}r_{i}$.
+The communication topology of the sensor network is described by a graph ${\mathcal{G}}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{1,2,\ldots,N\}$ denotes the sensor set, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the edge set defining the communication links among the sensors.
+For $i,j\in\mathcal{V}$, the edge $(j,i)$ indicates that sensor $j$ can transmit information to sensor $i$, and sensor $j$ is referred as a neighbor of sensor $i$.
+The neighbor set of sensor $i$ is denoted as $\mathcal{N}_{i}=\{j|(j,i)\in\mathcal{V}\}$, and $|\mathcal{N}_{i}|$ represents the number of neighbors of sensor $i$.
+The adjacency matrix is defined as $S=[a_{ij}]_{N\times N}$, where $a_{ij}=1$ if $(j,i)\in\mathcal{E}$ and $a_{ij}=0$ otherwise.
+Let $D=\text{diag}\{|\mathcal{N}_{1}|,\ldots,|\mathcal{N}_{N}|\}$, and the Laplacian matrix is defined as $\mathcal{L}=D-S$.
+An edge $(i,j)$ is undirected if $(i,j)\in\mathcal{E}$ implies $(j,i)\in\mathcal{E}$.
+The communication graph is termed undirected if all edges are undirected.
+The graph $G$ contains a directed path from sensor $i_{1}$ to sensor $i_{m}$, if there exists a sequence of connected edges $(i_{k},i_{k+1}),k=1,\ldots,m-1$.
+The communication graph is called strongly connected, if there exists a path between any pair of distinct nodes.
+Assumption 1
 $(C,A)$ is observable.
-2.3 Some Useful Lemmas
-Definition 1.
-[13] For a matrix $M\in\mathcal{R}^{n\times n}$, $M$ is said to be irreducibly diagonally dominant if
+Assumption 2
+The communication graph is directed and strongly connected.
+Remark 1
+Assumption 1 shows that observability can be achieved by the entire network, even if it is not satisfied by any single sensor.
+Assumption 2 can be extended to a jointly connected switching topology by applying the corresponding definition of communication networks [21, 14].
+II-C Problem Statement
   1. 1.
-    $M$ is irreducible,
+    Develop a distributed filter over directed graphs with a leader-following information fusion strategy using local measurements.
+Explore parameter design methods to guarantee the filter’s performance.
   2. 2.
-    $M$ is diagonally dominant, i.e., $|m_{ii}|\geq R^{{}^{\prime}}_{i}(M)$ for all $i=1,\ldots,n$,
-  3. 3.
-    There is an $i\in\{1,\ldots,n\}$ such that $|m_{ii}|>R^{{}^{\prime}}_{i}(M)$,
-where $R^{{}^{\prime}}_{i}(M)=\sum_{j\neq i}|m_{ij}|$.
-Lemma 1.
-[13] Let the matrix $M$ be irreducibly diagonally dominant. Then,
-  1. 1.
-    $M$ is nonsingular,
-  2. 2.
-    If $M$ is Hermitian and every main diagonal entry is positive, $M$ is positive definite.
-Lemma 2.
-[13] For a matrix $M$ and a positive integer $k$, it holds $\lim_{k\to\infty}M^{k}=0$ if and only if $\rho(M)<1$.
-Lemma 3.
-[13] For a nonnegative matrix $M=[m_{ij}]$, it holds
-$\text{min}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}\leq\rho(M)\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}.$ (2)
-Lemma 4.
-[26] For any matrix $M\in\mathcal{R}^{n\times n}$, it holds
-$\|M^{k}\|_{2}\leq\sqrt{n}\sum^{n-1}_{j=0}\binom{n-1}{j}\binom{k}{j}\|M\|^{j}_{2}\rho(M)^{k-j},$ (3)
-where $\binom{m}{n}$ is the combinatorial number with $\binom{k}{j}=0$ for $j>k$.
-2.4 Problem Statement
-  1. 1.
-    Develop a distributed filter algorithm over directed graphs, utilizing only measurements. Explore parameter design methods to guarantee the convergence and stability of the distributed filter.
-  2. 2.
-    Evaluate how the fusion step influences the steady-state and transient performance. Find out the performance gap between the proposed distributed filter and the centralized filter.
-3 Distributed Filter
-In this section, a consensus-on-only-measurement distributed filter (COMDF) is proposed.
-Subsequently, the state estimate error and the measurement estimate error are defined, and their properties are analyzed.
-Furthermore, two parameter design methods are provided to guarantee the stability and convergence of the proposed distributed filter.
-3.1 Design of the Distributed Filter
-This subsection proposes the consensus-on-only-measurement distributed filter, and an augmented leader-following measurement fusion strategy is designed to estimate the neighbors’ measurements.
-The state estimator structure of the target system (1) for sensor $i$ is designed as
-$\hat{x}_{i,k|k-1}=A\hat{x}_{i,k-1|k-1},{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}$ (4)
-$\hat{x}_{i,k|k}=\hat{x}_{i,k|k-1}+K(z^{(l)}_{i,k}-C\hat{x}_{i,k|k-1}),$ (5)
-where $z^{(l)}_{i,k}$ represents the sensor $i$’s estimate of measurements from other sensors at the $l$-th consensus step, the gain matrix $K$ is given by
-$K=PC^{T}(CPC^{T}+R)^{-1},$ (6)
-and $P$ is determined by solving the discrete algebraic Riccati equation
-$P=APA^{T}+Q-APC^{T}(CPC^{T}+R)^{-1}CPA^{T}.$ (7)
-The sensor $i$’s estimate of the sensor $j$’s measurement at the $l$-th fusion step, denoted as $z^{(l)}_{ij,k}$, is computed as
-$z^{(0)}_{ij,k}=C_{j}\hat{x}_{i,k|k-1},{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}$ (8)
-$z^{(l)}_{ij,k}$ $=z^{(l-1)}_{ij,k}-\mu_{ij}\Big{[}\sum^{N}_{l=1}a_{il}(z^{(l-1)}_{ij,k}-z^{(l-1)}_{lj,k})$ (9)
+    Find out the performance gap between the proposed distributed filter and the centralized filter. Evaluate how the fusion step number influences the performance.
+III Distributed Filter
+In this section, we propose a novel consensus-based distributed filter.
+Subsequently, we define the state estimation error and the measurement estimation error, and analyze their properties.
+Furthermore, we provide two parameter design methods to guarantee the filter’s performance.
+III-A Design of the Distributed Filter
+This subsection proposes a novel distributed state estimator over directed graphs with the consensus-based information fusion strategy.
+First, design sensor $i$’s estimate of sensor $j$’s measurement at the $l$-th fusion step, denoted as $z^{(l)}_{ij,k}$, as follows:
+$z^{(0)}_{ij,k}=C_{j}A\hat{x}_{i,k-1|k-1},$ (2)
+$z^{(l)}_{ij,k}$ $=z^{(l-1)}_{ij,k}-\mu_{ij}\Big[\sum_{t\in\mathcal{N}_{i}}a_{it}(z^{(l-1)}_{ij,k}-z^{(l-1)}_{tj,k})$ (3)
 $~~~~+a_{ij}(z^{(l-1)}_{ij,k}-y_{j,k})\Big],~~~l=1,2,\ldots,$
-where the consensus gain $\mu_{ij}$, designed later, is a positive constant, and $a_{ij}$ is the element of the adjacent matrix of the corresponding communication topology.
+where the consensus gain $\mu_{ij}$, to be designed later, is a positive constant, and $a_{ij}$ is the element of the adjacency matrix corresponding to the communication topology.
 By denoting the augmented vectors as $z^{(l)}_{i,k}=[(z^{(l)}_{i1,k})^{T},\ldots,(z^{(l)}_{iN,k})^{T}]^{T}$ and $y_{k}=[y^{T}_{1,k},\ldots,y^{T}_{N,k}]^{T}$, $z^{(l)}_{i,k}$ can be expressed as
-$z^{(0)}_{i,k}=C\hat{x}_{i,k|k-1},{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}$ (10)
-$~{}~{}~{}~{}z^{(l)}_{i,k}$ $=z^{(l-1)}_{i,k}-\Lambda_{i}\Big{[}\sum^{N}_{l=1}a_{il}(z^{(l-1)}_{i,k}-z^{(l-1)}_{l,k})$ (11)
+$z^{(0)}_{i,k}=CA\hat{x}_{i,k-1|k-1},$ (4)
+$~~~~z^{(l)}_{i,k}$ $=z^{(l-1)}_{i,k}-\Lambda_{i}\Big[\sum^{N}_{t=1}a_{it}(z^{(l-1)}_{i,k}-z^{(l-1)}_{t,k})~~~~$ (5)
 $~~~~+B_{i}(z^{(l-1)}_{i,k}-y_{k})\Big],$
 where $\Lambda_{i}=\text{diag}\{\mu_{i1}\otimes I_{r_{1}},\ldots,\mu_{iN}\otimes I_{r_{N}}\}$ and $B_{i}=\text{diag}\{a_{i1}\otimes I_{r_{1}},\ldots,a_{iN}\otimes I_{r_{N}}\}$.
-Now, the distributed estimator is constructed based on (4), (5), and (11).
-Remark 2.
-An augmented leader-following measurement information fusion strategy is designed in the algorithm, with only the transmission of measurement estimates in the sensor networks.
-The gain matrix $K$ can be precomputed, where the matrix $C$ and $R$ can be obtained by using a similar consensus algorithm as presented in (8) and (9).
-Two methods are presented here to obtain the corresponding matrices: one involves vectorizing the matrix and using the consensus strategy (9), while the other employs a matrix consensus strategy by replacing the vectors in (9) with the corresponding matrices.
-Remark 3.
-The proposed distributed estimator has the following advantages.
-First, it only needs the addition and substraction of the measurement vectors, thereby avoiding the computation and transmission of plenties of matrices.
-Second, the proposed distributed filter is asymptotically optimal as the fusion step $l$ tends to infinity.
-Third, it exhibits a stronger privacy protection performance with only the transmission of the measurements rather than the state estimates like [23, 14].
-Fourth, the algorithm can be employed in directed graphs, and the parameters can be designed by only using local neighbors’ information.
-Remark 4.
-The local gain matrix $K_{i}$, substituting for $K$ in (5), can also be utilized in the algorithm.
-By using local information regarding the measurement matrices and the measurement noise covariances, $K_{i}$ can be obtained and utilized.
-Under the local observability condition, this algorithm also works.
-To better present the properties of the proposed filter, the global $K$ is considered.
-There exist two parameters $\mu_{ij}$ and $l$, and the parameter design methods are introduced later.
-3.2 Two Estimation Errors
-Define the state estimation error and the measurement estimation error as
-$e_{i,k|k}=\hat{x}_{i,k|k}-x_{k},$
-and
-$\varepsilon^{(l)}_{i,k}=z^{(l)}_{i,k}-y_{k},~{}~{}~{}~{}$
-respectively.
-Denote the augmented vectors as $e_{k}=[e^{T}_{1,k},\ldots,e^{T}_{N,k}]^{T}$, $\varepsilon^{(l)}_{k}=[(\varepsilon^{(l)}_{1,k})^{T},\ldots,(\varepsilon^{(l)}_{N,k})^{T}]^{T}$, $z^{(l)}_{k}=[(z^{(l)}_{1,k})^{T},\ldots,(z^{(l)}_{N,k})^{T}]^{T}$, and $\nu_{k}=[\nu^{T}_{1,k},\ldots,\nu^{T}_{N,k}]^{T}$.
-Next, the statistical properties of the measurement estimation error $\varepsilon^{(l)}_{k}$ and the state estimation error $e_{i,k|k}$ are derived.
-Proposition 1.
-For the measurement estimation error $\varepsilon^{(l)}_{k}$ and its estimation error covariance $P_{\varepsilon,k}=E\{\varepsilon^{(l)}_{k}(\varepsilon^{(l)}_{k})^{T}\}$, the following results hold
+Next, the state estimator structure of the system (1) for sensor $i$ is designed as
+$\hat{x}_{i,k|k}=(A-KCA)\hat{x}_{i,k-1|k-1}+Kz^{(l)}_{i,k}$, (6)
+where the gain matrix $K$ is given by $K=\bar{P}C^{T}R^{-1},$ $\bar{P}=P-PC^{T}(CPC^{T}+R)^{-1}CP$, and $P$ is determined by solving the discrete algebraic Riccati equation $P=APA^{T}+Q-APC^{T}(CPC^{T}+R)^{-1}CPA^{T}$.
+Now, the distributed estimator is constructed based on (6), and its structure is illustrated in Fig. 1.
+Figure 1: Illustration of the proposed distributed estimator.
+Remark 2
+In the algorithm, both the measurement estimates $z^{(l-1)}_{it,k}$ and the raw measurements $y_{i,k}$ are transmitted.
+Between two consecutive time steps, the information is exchanged $l$ times.
+Since each node may not obtain sufficient measurements from its neighbors, the measurement estimates, obtained through prediction, compensate for this limitation.
+Remark 3
+The proposed distributed estimator offers several advantages.
+First, it complements existing distributed filters through a local measurement prediction fusion mechanism, thereby establishing a general framework for further investigation.
+Second, the proposed distributed filter demonstrates strong performance by achieving asymptotically optimality as the fusion step $l$ tends to infinity.
+Third, it can be applied to directed graphs with a simpler parameter design method compared to [5, 4].
+Fourth, when direct measurement information is unavailable, the predicted measurements compensate for missing data, thus maintaining estimation performance.
+Finally, its flexibility in communication steps makes it suitable for extensions to event-triggered or adaptive communication strategies, while the consensus framework can also be extended to support privacy-preserving and secure filtering schemes.
+III-B Two Estimation Errors
+This subsection computes two estimation errors for further analysis.
+Define the measurement estimation error and the state estimation error as $\varepsilon^{(l)}_{i,k}=z^{(l)}_{i,k}-y_{k},$ and $e_{i,k|k}=\hat{x}_{i,k|k}-x_{k},$ respectively.
+Denote the augmented vectors as
+$e_{k}=[e^{T}_{1,k},\ldots,e^{T}_{N,k}]^{T},~~~~~~~\varepsilon^{(l)}_{k}=[(\varepsilon^{(l)}_{1,k})^{T},\ldots,(\varepsilon^{(l)}_{N,k})^{T}]^{T},$
+$z^{(l)}_{k}=[(z^{(l)}_{1,k})^{T},\ldots,(z^{(l)}_{N,k})^{T}]^{T},~\nu_{k}=[\nu^{T}_{1,k},\ldots,\nu^{T}_{N,k}]^{T}.$
+Next, we derive the statistical properties of the measurement estimation error $\varepsilon^{(l)}_{k}$ and the state estimation error $e_{i,k|k}$.
+Additionally, we define their estimation error covariances as $P_{\varepsilon,k}=E\{\varepsilon^{(l)}_{k}(\varepsilon^{(l)}_{k})^{T}\}$ and $P_{k|k}=E\{e_{k}e^{T}_{k}\}$.
+Proposition 1
+The following statements hold
   1. 1.
-    The measurement estimation error $\varepsilon^{(l)}_{k}$ is given by
-    $\varepsilon^{(l)}_{k}=G^{l}\varepsilon^{(0)}_{k},$ (12)
-    where
-    $G$ $=I_{Nr}-\Lambda(\mathcal{L}\otimes I_{r}+B),~{}~{}~{}~{}~{}$ (13)
-    $\Lambda=\text{diag}\{\Lambda_{1},\ldots,\Lambda_{N}\},{}{}{}{}{}{}{}{}{}{}{}{}{}$ (14)
-    $B=\text{diag}\{B_{1},\ldots,B_{N}\},{}{}{}{}{}{}{}{}{}{}{}{}{}$ (15)
-    and
-    $\varepsilon^{(0)}_{k}$ $=(I_{N}\otimes CA)e_{k-1}-1_{N}\otimes\nu_{k}$ (16)
+    The measurement estimation error $\varepsilon^{(l)}_{k}$ is
+    $\varepsilon^{(l)}_{k}=G^{l}\varepsilon^{(0)}_{k}$,
+    where $G=I_{Nr}-\Lambda(\mathcal{L}\otimes I_{r}+B)$, $G^{l}$ represents the $l$-th power of $G$, $\Lambda=\text{diag}\{\Lambda_{1},\ldots,\Lambda_{N}\}$, $B=\text{diag}\{B_{1},\ldots,B_{N}\}$, and
+    $\varepsilon^{(0)}_{k}$ $=(I_{N}\otimes CA)e_{k-1}-1_{N}\otimes\nu_{k}$
     $~~~~-(I_{N}\otimes C)(1_{N}\otimes\omega_{k-1}).$
   2. 2.
-    The expectation value $E\{\varepsilon^{(l)}_{k}\}$ and $P_{\varepsilon,k}$ are given by
-    $E\{\varepsilon^{(l)}_{k}\}=G^{l}(I_{N}\otimes CA)E\{e_{k-1}\},$ (17)
-    and
+    The measurement estimation error covariance $P_{\varepsilon,k}$ is
     $P_{\varepsilon,k}$ $=G^{l}(I_{N}\otimes CA)P_{k-1}(I_{N}\otimes CA)^{T}(G^{l})^{T}$ (7)
-    $~{}~{}~{}~{}+G^{l}(I_{N}\otimes C)(U_{N}\otimes Q)(I_{N}\otimes C)^{T}(G^{l})^{T}$
-    $~{}~{}~{}~{}+G^{l}(U_{N}\otimes R)(G^{l})^{T},$
-    respectively.
-The proof of Proposition 1 is given in Appendix A.
-Lemma 5.
-Under the assumption that $\rho(G)<1$ , as $l$ tends to infinity, it holds
-$\lim_{l\to\infty}E\{\varepsilon^{(l)}_{k}\}=0,$ (19)
-and
-$\lim_{l\to\infty}P_{\varepsilon,k}=0.{}{}{}$ (20)
-The proof of Lemma 5 is given in Appendix B.
-Remark 5.
-Proposition 1 shows that $\varepsilon^{(l)}_{k}$ is a random variable.
-If $e_{k-1}$, $\omega_{k-1}$, and $\nu_{k}$ are Gaussian, $\varepsilon^{(l)}_{k}$ is Gaussian with the mean in (17) and the covariance $P_{\varepsilon,k}$ in (18).
-Lemma 5 demonstrates that as $l$ tends to infinity, $\varepsilon^{(l)}_{k}$ converges to 00.
-The state estimation error $e_{k}$ and the measurement estimation error $\varepsilon^{(l)}_{k}$ interact and influence each other.
-Proposition 2.
-The estimation error $e_{k}$ and the estimation error covariance $P_{k|k}=E\{e_{k}e^{T}_{k}\}$ are
-$e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}),$ (21)
-and
-$P_{k|k}$ $=\mathcal{A}(l)P_{k-1|k-1}\mathcal{A}^{T}(l)+\Phi(l),$ (22)
-respectively, where
-$\Phi(l)=\mathcal{B}(l)(U_{N}\otimes Q)\mathcal{B}^{T}(l)+\mathcal{D}(l)(U_{N}\otimes R)\mathcal{D}^{T}(l),$ (23)
-$\mathcal{A}(l)=I_{N}\otimes(A-KCA)+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$ (24)
+    $~~~~+G^{l}(I_{N}\otimes C)(U_{N}\otimes Q)(I_{N}\otimes C)^{T}(G^{l})^{T}$~~
+    $~~~~+G^{l}(U_{N}\otimes R)(G^{l})^{T}.$
+  3. 3.
+    The state estimation error $e_{k}$ is
+    $~~e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}),$
+    where
+    $~\mathcal{A}(l)=I_{N}\otimes(A-KCA)+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$ (8)
     $\mathcal{B}(l)=I_{N}\otimes(I-KC)-(I_{N}\otimes K)G^{l}(I_{N}\otimes C),$ (9)
@@
     $\mathcal{D}(l)=(I_{N}\otimes K)(I_{Nr}-G^{l}).$ (10)
-The proof of Proposition 2 is given in Appendix C.
-Remark 6.
-Based on Proposition 2, two parameters still need to be designed: one is the parameter $\mu_{ij}$ in $G$, and the other is the fusion step $l$.
-Observing (24), $\rho(G)<1$ ensures the convergence of the filter.
-Subsequently, we focus on the design of the parameters $\mu_{ij}$ and $l$.
-3.3 Parameter Design for $\mu_{i}$
-The objective of designing the parameter $\mu_{i}$ is to ensure that $\rho(G)<1$, and it can guarantee the convergence of the measurement estimate error and the state estimate error.
-Next, two methods are provided for the design of the parameter $\mu_{i}$.
-3.3.1 Distributed Design
-A distributed design approach for $\mu_{ij}$ is proposed to circumvent the utilization of the global topology information for each sensor.
-First, the elements of $G$ are presented, laying the foundation for subsequent design considerations.
-Define the submatrix $G_{[ij]}$ as
-$G_{[ij]}$ $=[0_{r\times r(i-1)},I_{r},0_{r\times r(N-i)}]G$
-×$~{}~{}~{}~{}\times[0^{T}_{r\times r(j-1)},I^{T}_{r},0^{T}_{r\times r(N-j)}]^{T},$
-where $[ij]$ represents the $i$-th row and $j$-th column index of the corresponding matrix block. The detailed submatrices are
-$G_{[ii]}$ $=\text{diag}\{g_{ii,1},\dots,g_{ii,N}\},$
-and
-$G_{[ij]}$ $=\text{diag}\{g_{ij,1},\dots,g_{ij,N}\},$
+  4. 4.
+    The state estimation error covariance $P_{k|k}$ is
+    $~~P_{k|k}=\mathcal{A}(l)P_{k-1|k-1}\mathcal{A}^{T}(l)+\Phi(l),$ (11)
     where
-$~{}~{}~{}~{}~{}g_{[ii,l]}$ $=(1-\mu_{il}(l_{ii}+a_{il}))\otimes I_{r_{l}},$ (27)
-and
-$g_{[ij,l]}$ $=(-\mu_{il}l_{ij})\otimes I_{r_{l}}.~{}~{}~{}~{}~{}~{}~{}$ (28)
-Consider the row sum of the block matrix, and it holds
-$\sum^{n}_{j=1}g_{[ij,l]}$ $=\Big{(}1-\mu_{il}\sum^{N}_{j=1}{l_{ij}}-\mu_{il}a_{il}\Big{)}\otimes I_{r_{l}}$ (29)
-$=\Big{(}1-\mu_{il}a_{il}\Big{)}\otimes I_{r_{l}}.$
-Next, a reasonable range of the parameter $\mu_{ij}$ is provided to make $\rho(G)<1$.
-Lemma 6.
-Under Assumption 1, if $\mu_{il}$ satisfies
-$0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}},$
-then it holds $\rho(G)<1$.
-The proof of Lemma 6 is given in Appendix D.
-Remark 7.
-The distributed design method can be employed to design parameter $\mu_{ij}$ by utilizing the local information instead of relying on global topology information [10, 3].
-Additionally, It is suggested to design $\mu_{ij}$ such that $1-\mu_{il}(l_{ii}+a_{il})$ approaches 00, leading to a smaller $\rho(G)$ with a higher probability.
-3.3.2 Unified Design
-A unified design method is also provided for undirected graphs. Consider the scenario where all $\mu_{i}$ are identical, denoted as $\mu$. Subsequently, the symbol $G$ is redefined as follows:
-$G$ $=I_{Nr}-\mu(\mathcal{L}\otimes I_{r}+B).$
-Based on Lemma 1, $G$ is positive definite.
+    $\Phi(l)=\mathcal{B}(l)(U_{N}\otimes Q)\mathcal{B}^{T}(l)+\mathcal{D}(l)(U_{N}\otimes R)\mathcal{D}^{T}(l).$ (12)
+The proof of Proposition 1 is given in Appendix A.
+Remark 4
+Proposition 1 shows that $\varepsilon^{(l)}_{k}$ is a random variable with the mean $E\{\varepsilon^{(l)}_{k}\}=G^{l}(I_{N}\otimes CA)E\{e_{k-1}\}$ and the covariance $P_{\varepsilon,k}$.
+Under the assumption that $\rho(G)<1$, as $l$ tends to infinity, it holds $\lim_{l\to\infty}E\{\varepsilon^{(l)}_{k}\}=0$ and $\lim_{l\to\infty}P_{\varepsilon,k}=0$.
+The state estimation error $e_{k}$ and the measurement estimation error $\varepsilon^{(l)}_{k}$ interact and influence each other.
+To improve the filter performance, two parameters will be designed: one is the parameter $\mu_{ij}$ in $G$, and the other is the fusion step $l$.
+III-C Parameter Design for $\mu_{i}$
+This subsection will design the parameter $\mu_{i}$ to ensure the convergence of two estimation errors.
+III-C1 Distributed Design
+We propose a distributed design approach for $\mu_{ij}$ to circumvent the utilization of the global topology information for each sensor.
+First, define the submatrix $G_{[ij]}$ as $G_{[ij]}=[0_{r\times r(i-1)},I_{r},0_{r\times r(N-i)}]G[0^{T}_{r\times r(j-1)},I^{T}_{r},0^{T}_{r\times r(N-j)}]^{T},$ where $[ij]$ represents the $i$-th row and $j$-th column index of the corresponding matrix block.
+The detailed submatrices are $G_{[ii]}=\text{diag}\{g_{ii,1},\dots,g_{ii,N}\}$ and $G_{[ij]}=\text{diag}\{g_{ij,1},\dots,g_{ij,N}\}$, where
+$~~~~g_{[ii,l]}$ $=(1-\mu_{il}(l_{ii}+a_{il}))\otimes I_{r_{l}},$ (13)
+$g_{[ij,l]}$ $=(-\mu_{il}l_{ij})\otimes I_{r_{l}},$~~~~~~~ (14)
+and $l_{ij}$ is the $i$-th row and $j$-th column element of the Laplacian matrix $\mathcal{L}$. Next, we introduce a method to select the parameter $\mu_{ij}$ from a reasonable range, given by:
+$0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}}.$~~~~~~~~~~~~ (15)
+We provide the following lemma to illustrate the choice of $\mu_{il}$ satisfying (15) can ensure $\rho(G)<1$.
+Lemma 5
+Under Assumption 2, if $\mu_{il}$ satisfies (15) then it holds $\rho(G)<1$.
+The proof of Lemma 5 is given in Appendix B.
+III-C2 Unified Design
+A unified design method is also provided for undirected graphs.
+Consider the scenario where all $\mu_{i}$ are identical, denoted as $\mu$.
+The symbol $G$ is then redefined as $G=I_{Nr}-\mu(\mathcal{L}\otimes I_{r}+B).$ Based on Lemma 1, $G$ is positive definite.
 By selecting $\mu<\frac{1}{\rho(\mathcal{L}\otimes I_{r}+B)}$, it is ensured that $\rho(G)<1$.
-This method needs the global communication topology information.
-3.4 Parameter Design for $l$
-This subsection aims to design the parameter $l$ to ensure the convergence of the distributed estimators.
-By employing Lemma 2 and considering (24), it is evident that $\rho(\mathcal{A}(l))<1$ can guarantee the convergence of the distributed filter.
-The following results provide a lower bound of the consensus step $l$.
-Theorem 1.
-Under Assumption 1 and 2, the estimation error covariance $P_{i,k|k}$ is uniformly upper-bounded, if $\|G\|_{2}<1$ and $l>l_{0}$, where
-$l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}.$ (30)
-The proof of Theorem 1 is given in Appendix E.
-Remark 8.
-For Theorem 1, there are two elements that decide the value of $l_{0}$, i.e., $\|G\|_{2}$ and $\bar{K}=\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}$.
-On one hand, the term $\|G\|_{2}$ is relevant to the communication graph and the parameter $\mu_{i}$.
-The smaller $\|G\|_{2}$, the smaller $l_{0}$.
+This method requires the global communication topology information.
+III-D Parameter Design for $l$
+This subsection aims to provide the lower bound of the fusion step number $l$ to ensure the convergence of the distributed estimators.
+Theorem 1
+Under Assumptions 1 and 2, if $\|G\|_{2}<1$ and $l>l_{0}$, where
+$~l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}},$ (16)
+then $\mathcal{A}(l)$ is Schur stable, and the estimation error covariance $P_{i,k|k}$ is uniformly upper-bounded.
+The proof of Theorem 1 is given in Appendix C.
+Remark 5
+From Theorem 1, the value $l_{0}$ is determined by two factors: $\|G\|_{2}$ and $\bar{K}=\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}$.
+On one hand, the term $\|G\|_{2}$ is related to the communication graph and the parameter $\mu_{i}$.
+The smaller $\|G\|_{2}$ is , the smaller $l_{0}$ becomes.
 On the other hand, the term $\bar{K}$ indicates the stability margin of the estimator.
-The larger $\bar{K}$, the smaller $l_{0}$.
-4 Performance Analysis
-In this section, the centralized estimator is first introduced.
-Then, the steady-state performance gap between the centralized filter and the proposed distributed filter is analyzed with the increasing fusion step $l$.
-Additionally, the transient performance is also taken into consideration.
-4.1 Centralized Estimator
-Consider a centralized estimator as follows:
-$\hat{x}_{c,k|k-1}=A\hat{x}_{c,k-1|k-1},$ (31)
-$\hat{x}_{c,k|k}=\hat{x}_{c,k|k-1}+K(y_{k}-C\hat{x}_{c,k|k-1}),$ (32)
-where $y_{k}$, $K$, and $C$ are given in (6).
-For the centralized estimator, denote the estimation error as $e_{c,k|k}=\hat{x}_{c,k|k}-x_{k}$, the corresponding estimation error covariance as $P_{c,k|k}=E\{e_{c,k|k}e^{T}_{c,k|k}\}$, the augmented estimation error as $e_{cc,k}=1_{N}\otimes e_{c,k}$, and the augmented estimation error covariance as $P_{cc,k|k}=E\{e_{cc,k|k}e^{T}_{cc,k|k}\}=U_{N}\otimes P_{c,k|k}.$
-By combining the estimator (31), (32) and the dynamical system (1), the estimation error $e_{c,k}$ is
-$e_{c,k}$ $=(A-KCA)\hat{x}_{c,k-1|k-1}+Ky_{k}-Ax_{k-1}-\omega_{k-1}$ (33)
-$=(A-KCA)e_{c,k-1}-(I-KC)\omega_{k-1}+K\nu_{k}.$
-Then, $P_{c,k|k}=E\{e_{c,k|k}e^{T}_{c,k|k}\}$ can be computed as
-$P_{c,k|k}$ $=(A-KCA)P_{c,k-1|k-1}(A-KCA)^{T}$ (34)
-$~{}~{}~{}+(I-KC)Q(I-KC)^{T}+KRK^{T}.$
-Similarly, the augmented estimation error of the centralized estimator $e_{cc,k}=1_{N}\otimes e_{c,k}$ and the corresponding estimation error covariance $P_{cc,k|k}$ can be obtained as
-$e_{cc,k}=\mathcal{A}_{cc}e_{cc,k-1}-\mathcal{B}_{cc}(1_{N}\otimes\omega_{k-1})+\mathcal{D}_{cc}(1_{N}\otimes\nu_{k}),$ (35)
-and
-$P_{cc,k|k}$ $=\mathcal{A}_{cc}P_{cc,k-1|k-1}\mathcal{A}^{T}_{cc}+\Phi_{cc},~{}~{}~{}~{}~{}~{}~{}~{}$ (36)
-respectively, where
-$\Phi_{cc}=\mathcal{B}_{cc}(U_{N}\otimes Q)\mathcal{B}^{T}_{cc}+\mathcal{D}_{cc}(U_{N}\otimes R)\mathcal{D}^{T}_{cc},$ (37)
+The larger $\bar{K}$ is, the smaller $l_{0}$ becomes.
+IV Performance Analysis
+This section analyzes the performance of the proposed distributed filter.
+Specifically, we present the steady-state performance gap between the centralized filter and the proposed distributed filter as the fusion step number $l$ increases.
+IV-A Centralized Estimator
+Consider a centralized estimator for comparison, given by:
+$\hat{x}_{c,k|k}=A\hat{x}_{c,k-1|k-1}+K(y_{k}-CA\hat{x}_{c,k-1|k-1}),$ (17)
+Denote the estimation error as $e_{c,k|k}=\hat{x}_{c,k|k}-x_{k}$, the estimation error covariance as $P_{c,k|k}=E\{e_{c,k|k}e^{T}_{c,k|k}\}$, the augmented estimation error as $e_{cc,k}=1_{N}\otimes e_{c,k}$, and the augmented estimation error covariance as $P_{cc,k|k}=E\{e_{cc,k|k}e^{T}_{cc,k|k}\}=U_{N}\otimes P_{c,k|k}.$ By combining the estimator (17) and the dynamical system (1), $e_{cc,k}=1_{N}\otimes e_{c,k}$ can be computed as
+$~~~e_{cc,k}=\mathcal{A}_{cc}e_{cc,k-1}-\mathcal{B}_{cc}(1_{N}\otimes\omega_{k-1})+\mathcal{D}_{cc}(1_{N}\otimes\nu_{k}),$ (18)
+where
 $\mathcal{A}_{cc}=I_{N}\otimes(A-KCA),$ (19)
@@
 $\mathcal{D}_{cc}=I_{N}\otimes K.$ (21)
-Remark 9.
-This centralized estimator utilizes the fixed gain matrix $K$, resulting in a suboptimal filter at every step.
-Nevertheless, as the time step $k$ approaches infinity, $P_{c,k|k}$ converges to $P$ in (7).
-Therefore, the relations between the steady-state performance of $P_{i,k|k}$ and $P_{c,k|k}$ reflect those between the steady-state performance of $P_{i,k|k}$ and $P$.
-4.2 Convergence Analysis
-This section investigates the steady-state performance gap between $P_{cc,k|k}$ and $P_{k|k}$, and sheds light on the influence of the fusion step $l$ on the performance.
-Section 3.3 has shown that the parameter $\mu_{ij}$ can be designed to ensure that $\rho(G)<1$.
+The corresponding estimation error covariance $P_{cc,k|k}$ is
+$P_{cc,k|k}$ $=\mathcal{A}_{cc}P_{cc,k-1|k-1}\mathcal{A}^{T}_{cc}+\Phi_{cc},$~~~~~~~~~~ (22)
+where
+$\Phi_{cc}=\mathcal{B}_{cc}(U_{N}\otimes Q)\mathcal{B}^{T}_{cc}+\mathcal{D}_{cc}(U_{N}\otimes R)\mathcal{D}^{T}_{cc}.$ (23)
+Since we analyze the steady-state performance gap, the centralized filter with the fixed gain will converge to the optimal solution.
+IV-B Convergence Analysis
+This section investigates the steady-state performance gap between $P_{cc,k|k}$ and $P_{k|k}$, and sheds light on the influence of the fusion step number $l$ on the performance.
+Section III-C has shown that the parameter $\mu_{ij}$ can be designed to ensure that $\rho(G)<1$.
 Therefore, we make the following assumption.
-Assumption 3.
+Assumption 3
 The matrix $G$ satisfies $\rho(G)<1$, and $\mathcal{A}(l)$ is Schur stable.
-Theorem 2.
-Under Assumption 1, 2, and 3, $P_{k|k}$ and $P_{cc,k|k}$ converge to the unique solutions of the discrete-time Lyapunov equations (DLE)
-$P^{(l)}$ $=\mathcal{A}(l)P^{(l)}\mathcal{A}^{T}(l)+\Phi(l),$ (41)
+Remark 6
+Assumption 3 can be satisfied through the proposed parameter design method, as established in Lemma 5 and Theorem 1.
+In other words, this assumption indicates that the parameters have been appropriately designed.
+Theorem 2
+Under Assumptions 1, 2, and 3, $P_{k|k}$ and $P_{cc,k|k}$ converge to the unique solutions of the discrete-time Lyapunov equations (DLE)
+$P^{(l)}$ $=\mathcal{A}(l)P^{(l)}\mathcal{A}^{T}(l)+\Phi(l),$~~ (24)
 and
-$P_{cc}$ $=\mathcal{A}_{cc}P_{cc}\mathcal{A}^{T}_{cc}+\Phi_{cc},$ (42)
+$P_{cc}$ $=\mathcal{A}_{cc}P_{cc}\mathcal{A}^{T}_{cc}+\Phi_{cc},$~~~~~~~~ (25)
 respectively, where $\mathcal{A}(l)$, $\Phi(l)$, $\mathcal{A}_{cc}$, and $\Phi_{cc}$ have been defined in (8), (12), (19), and (23), respectively.
-The proof of Theorem 2 is given in Appendix F.
-To evaluate the effect of the fusion step $l$ on the steady-state performance of the proposed filter, we introduce the following notations
-$\bar{\mathcal{A}}(l)=\mathcal{A}(l)-\mathcal{A}_{cc}=(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$
+The proof of Theorem 2 is given in Appendix D.
+To evaluate the effect of the fusion step number $l$ on the steady-state performance, we introduce the following notations
+$\bar{\mathcal{A}}(l)=\mathcal{A}(l)-\mathcal{A}_{cc}=(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$ (26)
 […]
-and
 $\mathcal{\bar{D}}(l)=\mathcal{D}(l)-\mathcal{D}_{cc}=-(I_{N}\otimes K)G^{l}.$
-Lemma 7.
+Lemma 6
 The difference between $P^{(l)}$ and $P_{cc}$ is
@@
 $~~~+\mathcal{D}_{cc}(U_{N}\otimes R){\mathcal{\bar{D}}}^{T}(l)+{\mathcal{\bar{D}}}(l)(U_{N}\otimes R){\mathcal{\bar{D}}}^{T}(l).$
-The proof of Lemma 7 is given in Appendix G.
-Theorem 3.
-Under Assumption 1 and 2, there exist constants $M_{1},M_{2}>0$, such that
+The proof of Lemma 6 is given in Appendix E.
+Theorem 3
+Under Assumptions 1 and 2, there exist constants $M_{1},M_{2}>0$, such that
   1. 1.
@@
     $\|P^{(l)}-P_{cc}\|_{2}\leq M_{2}\|G\|^{l}_{2}.$ (30)
-The proof of Theorem 3 is given in Appendix H.
-Remark 10.
-Two conditions related to the communication graph are presented, since the mere satisfaction of a spectral radius of the matrix less than 1 does not ensure the spectral norm of the matrix less than 1.
-For an undirected graph, the spectral radius $\rho(G)$ is equal to the spectral norm $\|G\|_{2}$, and the conditions of item (1) and (2) in Theorem 3 are identical.
-For both items, as the fusion step $l$ tends to infinity, $\|P^{(l)}-P_{cc}\|_{2}$ converges to 00 according to (45) and (46).
-4.3 Transient Performance
-The transient performance of the proposed filter is analyzed in the following.
-Lemma 8.
-Under Assumption 3, as $l$ tends to infinity, it holds that
-$\lim_{l\to\infty}\mathcal{A}(l)=\mathcal{A}_{cc},$
-$\lim_{l\to\infty}\mathcal{B}(l)=\mathcal{B}_{cc},$
-and
-$\lim_{l\to\infty}\mathcal{D}(l)=\mathcal{D}_{cc}.$
-The proof of Lemma 8 is given in Appendix I.
-Theorem 4.
-Under Assumption 1, 2, and 3, and considering the case that $P_{k-1|k-1}=P_{cc,k-1|k-1}$, one has
-$\lim_{l\to+\infty}P_{i,k|k}=P_{c,k|k}.$ (47)
-The proof of Theorem 4 is given in Appendix J.
-Remark 11.
-Based on the structural characteristics of $\mathcal{A}(l)$, $\mathcal{B}(l)$, and $\mathcal{D}(l)$, Lemma 8 shows that the matrices $\mathcal{A}(l)$, $\mathcal{B}(l)$, and $\mathcal{D}(l)$ approach to the centralized matrices $\mathcal{A}_{cc}$, $\mathcal{B}_{cc}$, and $\mathcal{D}_{cc}$, respectively, as the fusion step tends to infinity.
-Then, it is demonstrated that the transient performance exhibits similar results.
-In other words, as the fusion step increases, the performance of the proposed filter approaches that of the centralized filter.
-By considering both the transient properties and the convergence properties, we can gain comprehensive insights into the performance of the proposed algorithm.
-5 Simulations
+The proof of Theorem 3 is given in Appendix F.
+Remark 7
+Two conditions related to the communication graph are presented, since ensuring that the spectral radius of the matrix is less than 1 does not necessarily imply that its spectral norm is also less than 1.
+For an undirected graph, the spectral radius $\rho(G)$ equals the spectral norm $\|G\|_{2}$, and the conditions in item (1) and (2) of Theorem 3 are identical.
+In both cases, as the fusion step number $l$ approaches infinity, $\|P^{(l)}-P_{cc}\|_{2}$ converges to 0, as demonstrated in (29) and (30).
+V Simulations
 In this section, the effectiveness of the theoretical results is validated through a target tracking numerical experiment.
 Consider a sensor network comprising five sensors labeled from 1 to 5, and its communication topology is illustrated in Fig.
-1.
+2.
 Figure 2: The diagram of the communication topology.
-Consider the target dynamics described by
-$A=\left(\begin{array}[]{cc}I_{2}&TI_{2}\\ 0&I_{2}\end{array}\right),$ (48)
-where $T=0.25$ is the discretization interval. The process noise covariance is defined as
+Consider the target dynamics described by $A=\text{diag}\{I_{2}~TI_{2};0~I_{2}\}$, where $T=0.25$ is the discretization interval. The process noise covariance is defined as
 $\bar{Q}=\left(\begin{array}[]{cc}\frac{T^{3}}{3}&\frac{T^{2}}{2}\\ \frac{T^{2}}{2}&T\end{array}\right),~~~Q=\left(\begin{array}[]{cc}\bar{Q}&0.5\bar{Q}\\ 0.5\bar{Q}&\bar{Q}\end{array}\right).$ (31)
-Two kinds of sensors are employed in the sensor network: position sensors and velocity sensors. The observation matrix for the position sensors is given by
-$C_{p}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\end{array}\right),$ (50)
-with the measurement noise covariance $R_{p}=\text{diag}\{1,1\}$. Similarly, the observation matrix for the velocity sensors is represented by
-$C_{v}=\left(\begin{array}[]{cccc}0&0&1&0\\ 0&0&0&1\end{array}\right),$ (51)
-with the measurement noise covariance $R_{v}=\text{diag}\{5,5\}$.
-In this sensor network, it is assumed that sensor 1, 2, and 4 are the position sensors, while sensor 3 and 5 serve as the velocity sensors.
+Two kinds of sensors are employed in the sensor network: position sensors and velocity sensors.
+The observation matrix for the position sensors is given by $C_{p}=[I_{2}~0_{2}]$ with the measurement noise covariance $R_{p}=\text{diag}\{1,1\}$.
+Similarly, the observation matrix for the velocity sensors is represented by $C_{v}=[0_{2}~I_{2}]$ with the measurement noise covariance $R_{v}=\text{diag}\{5,5\}$.
+In this sensor network, sensors 1, 2, and 4 are designated as the position sensors, while sensors 3 and 5 serve as the velocity sensors.
 The initial state is set as $x_{0}=[1;1;1;1]$, and the initial state estimate is set as a random variable with the mean $\hat{x}_{i,k|k}=x_{0}$ and the initial estimation covariance $P_{i,0|0}=\text{diag}\{10,10,10,10\}$.
@@
 The mean square error (MSE) is utilized to evaluate the performance of the estimator based on the Monte Carlo method, described by
-$MSE_{i,k}=\frac{1}{M}\sum^{M}_{l=1}\|\hat{x}^{(l)}_{i,k}-x^{(l)}_{k}\|^{2}_{2},$ (52)
-where $M$ is the trial number, and $\hat{x}^{(l)}_{i,k}$ and $x^{(l)}_{k}$ are the state estimate and the true state at the $l$-th trial, respectively.
+$MSE_{i,k}=\frac{1}{M}\sum^{M}_{m=1}\|\hat{x}^{(m)}_{i,k}-x^{(m)}_{k}\|^{2}_{2},$ (32)
+where $M$ is the trial number, and $\hat{x}^{(m)}_{i,k}$ and $x^{(m)}_{k}$ are the state estimate and the true state at the $m$-th trial, respectively.
 In the simulation, $M$ is set as $1000$.
-All simulations are executed in MATLAB R2020a with an Intel Core i5-1135G7 CPU @ 2.40 GHz.
+All simulations are executed in MATLAB R2024b with an Intel Core Ultra 9 185H CPU @ 2.30 GHz.
 Two examples are designed to verify the effectiveness of the proposed algorithm.
-Example 1 aims to demonstrate the steady-state performance of the proposed COMDF as the fusion step $l$ increases.
-Example 2 is designed to show the performances and properties of COMDF in comparison with other existing consensus-based distributed filers.
-Figure 2: Illustration figure for the steady-state performance of COMDF with the increasing fusion step $l$.
+We refer to our proposed consensus-based distributed filter as CBDF.
+Example 1 aims to demonstrate the steady-state performance of the proposed filter as the fusion step number $l$ increases.
+Example 2 is designed to show the performances and properties in comparison with other existing distributed filers.
+Figure 3: Illustration figure for the steady-state performance of CBDF with the increasing fusion step number $l$.
 Example 1: Fig.
-2 exhibits the steady-state performance of COMDF of five sensors with the increasing fusion step $l$.
-It is shown that the performance gap between the centralized filter and COMDF is exponential convergence (Theorem 3).
-In addition, a small fusion step can also ensure the performance of COMDF.
-Example 2: To assess the performance of the proposed COMDF, three other algorithms are considered: the consensus-on-measurement distributed filter (CMDF) from [6], the consensus-on-information distributed filter (CIDF) from [4], and the centralized Kalman filter (CKF) in Section 4.1.
+3 exhibits the steady-state performance of CBDF of five sensors with the increasing fusion step number $l$.
+It is shown that the performance gap between the centralized filter and CBDF is exponential convergence (Theorem 3).
+In addition, a small fusion step can also ensure the performance of CBDF.
+Figure 4: The steady-state error of sensor 1 across 1000 Monte Carlo simulations for four algorithms.
+In each box plot, the red dotted line represents the mean, while the blue solid line indicates the median.
+The upper and lower edges of each box correspond to the 25th and 75th percentiles, respectively.
+Red circles are used to highlight outliers that fall beyond 1.5 times the interquartile range.
+Example 2: To assess the performance of the proposed CBDF, three other algorithms are considered: the consensus-on-information distributed filter (CIDF) from [5], asymptotically optimal distributed filter (CMDF) from [4], and the centralized Kalman filter (CKF) in Section IV-A.
 For comparison, it is assumed that the communication topology is undirected in Fig.
-1, and the fusion step is set as $l=10$.
+2, and the fusion step is set as $l=10$.
+The Metropolis weights are adopted in CIDF and CMDF.
+We define the steady-state error for sensor $i$ as
+$E_{i,\infty}=\|\hat{x}_{i,\infty}-x_{\infty}\|^{2}_{2}.$ (33)
 Fig.
-3 displays the performance of four algorithms with the increasing time step $k$.
-As the time step $k$ tends to infinity, four algorithms converge.
-CKF and COMDF are the algorithms that fix the gain matrix $K$, hence, they may not be optimal in every step.
-However, as $k$ tends to infinity, CKF approaches optimality.
-Meanwhile, the performance of COMDF approaches that of CKF with the increasing fusion step.
-Consequently, before the algorithms converge, CIDF and CMDF show a faster convergence velocity.
-CIDF adopts the covariance intersection method to handle the noise correlations, and the steady-state performance is degraded.
-Figure 3: Illustration figure for the performance of four algorithms with the increasing time step $k$.
-Table 1 presents the time consumption and the memory usage of four algorithms in Example 2.
+4 shows the steady-state error of sensor 1 for four algorithms from 1000 Monte Carlo experiments.
+It is observed that CKF exhibits the best performance among the four algorithms, while CBDF demonstrates good performance.
+It is worth mentioning that this simulation is executed over undirected graphs for comparison, since the weight design methods are not well established for directed graphs [20].
+Table I presents the execution time and memory usage of the four algorithms in Example 2.
 All four algorithms exhibit similar memory usage, but substantial differences emerge in terms of time consumption.
-Notably, COMDF demonstrates the smallest time consumption among the three distributed algorithms, attributed to its transmission of only measurements.
-Conversely, CIDF exhibits the highest time consumption among the three distributed filters, possibly due to a higher frequency of the inverse operation in the algorithm.
-Table 1: The time consumption and the memory usage of four algorithms in Example 2.
-Algorithm Time Consumption (s) Memory Usage (MB)
-COMDF      23.864      2192
-CMDF      32.709      2183
-CIDF      144.438      2189
-CKF      2.005      2194
-6 Conclusions
-This paper proposes a consensus-on-measurement distributed filter over directed graphs, embedded with an augmented leader-following measurement fusion strategy.
-The parameters are designed to guarantee the uniformly upper bound of the estimation error covariances.
-The steady-state and transient performances are analyzed with the increasing fusion step, and the relations between the proposed distributed filter and the centralized filter are revealed.
-In the future, it is desired to design a dynamic gain matrix to optimize the distributed filter at every step, aiming for a faster convergence.
-Additionally, there is an intention to explore a new parameter design method to reduce the spectral radius of the communication matrix.
+The CBDF requires less computation time, owing to its reliance primarily on simple addition and subtraction operations.
+TABLE I: The execution time and memory usage in Example 2.
+Algorithm Execution Time (s) Memory Usage (MB)
+CBDF 5.21 3294
+CMDF 14.49 3341
+CIDF 41.69 3295
+CKF 1.07 3326
+VI Conclusion
+This paper proposed a consensus-based distributed filter over directed graphs, embedded with an augmented leader-following information fusion strategy.
+The parameters were designed to guarantee the uniform upper bound on the estimation error covariances.
+The steady-state performance was analyzed as the fusion step number increases, and the relations between the proposed distributed filter and the centralized filter were revealed.
+In the future, we aim to design a dynamic gain matrix to optimize the distributed filter at each step, with the goal of achieving faster convergence.
+Additionally, we intend to explore a new parameter design method to reduce the spectral radius of the communication matrix.
 Appendix A PROOF of Proposition 1
 Item 1): Using equation (5), $\varepsilon^{(l)}_{i,k}$ can be derived as
-$\varepsilon^{(l)}_{i,k}$ $=z^{(l-1)}_{i,k}-\Lambda_{i}\Big{[}\sum^{N}_{l=1}a_{il}(z^{(l-1)}_{i,k}-z^{(l-1)}_{l,k})$ (53)
+$\varepsilon^{(l)}_{i,k}$ $=z^{(l-1)}_{i,k}-\Lambda_{i}\Big[\sum^{N}_{l=1}a_{il}(z^{(l-1)}_{i,k}-z^{(l-1)}_{l,k})$
 $~~~~+B_{i}(z^{(l-1)}_{i,k}-y_{k})\Big]-y_{k}$
 $=\varepsilon^{(l-1)}_{i,k}-\Lambda_{i}\Big[\sum^{N}_{l=1}a_{il}(\varepsilon^{(l-1)}_{i,k}-\varepsilon^{(l-1)}_{l,k})+B_{i}\varepsilon^{(l-1)}_{i,k}\Big$].
-Then, the augmented vector $\varepsilon^{(l)}_{k}$ can be obtained as
-$\varepsilon^{(l)}_{k}$ $=G\varepsilon^{(l-1)}_{k},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (54)
-where $G$ is given in (13). When $l=0$, it follows
-$\varepsilon^{(0)}_{i,k}$ $=z^{(0)}_{i,k}-y_{k}$ (55)
-$=CAe_{i,k-1}-C\omega_{k-1}-\nu_{k}.$
-Similarly, the augmented vector $\varepsilon^{(0)}_{k}$ is
-$\varepsilon^{(0)}_{k}$ $=(I_{N}\otimes CA)e_{k-1}-(I_{N}\otimes C)(1_{N}\otimes\omega_{k-1})$ (56)
-$~{}~{}~{}~{}-1_{N}\otimes\nu_{k}.$
-Next, it can be concluded that
-$\varepsilon^{(l)}_{k}$ $=G\varepsilon^{(l-1)}_{k}$ (57)
-$=G^{l}\varepsilon^{(0)}_{k}.$
+Then, the augmented vector $\varepsilon^{(l)}_{k}$ can be obtained as $\varepsilon^{(l)}_{k}=G\varepsilon^{(l-1)}_{k}$.
+When $l=0$, it follows $\varepsilon^{(0)}_{i,k}=z^{(0)}_{i,k}-y_{k}=CAe_{i,k-1}-C\omega_{k-1}-\nu_{k}.$ Similarly, the augmented vector $\varepsilon^{(0)}_{k}$ is $\varepsilon^{(0)}_{k}=(I_{N}\otimes CA)e_{k-1}-(I_{N}\otimes C)(1_{N}\otimes\omega_{k-1})-1_{N}\otimes\nu_{k}.$ Next, it can be concluded that $\varepsilon^{(l)}_{k}=G\varepsilon^{(l-1)}_{k}=G^{l}\varepsilon^{(0)}_{k}$.
 Item 2): Based on the results of Item 1), $E\{\varepsilon^{(l)}_{k}\}$ and $E\{\varepsilon^{(l)}_{k}(\varepsilon^{(l)}_{k})^{T}\}$ can be calculated.
-Appendix B PROOF of Lemma 2
-According to Lemma 2, since $\rho(G)<1$, it follows $\lim_{l\to+\infty}G^{l}=0$. Observing (17) and (18), $G^{l}$ exists in all terms. Hence, this lemma can be proven.
-Appendix C PROOF of Proposition 2
-Utilizing the state estimator equations (4) and (5) in conjunction with the dynamical system (1), the estimation error $e_{i,k}$ for sensor $i$ is
+Item 3): Utilizing the state estimator (6) in conjunction with the dynamical system (1), the estimation error $e_{i,k}$ for sensor $i$ is
 $e_{i,k}$ $=(A-KCA)\hat{x}_{i,k-1|k-1}+Kz^{(l)}_{i,k}-Ax_{k-1}-\omega_{k-1}$
-$=(A-KCA)e_{i,k-1}+K\varepsilon^{(l)}_{i,k}$
-$~{}~{}~{}~{}-(I-KC)\omega_{k-1}+K\nu_{k}.$
+$=(A-KCA)e_{i,k-1}+K\varepsilon^{(l)}_{i,k}-(I-KC)\omega_{k-1}+K\nu_{k}.$
 Subsequently, by using Proposition 1, the augmented estimation error $e_{k}$, i.e., $e_{k}=[e^{T}_{1,k},\ldots,e^{T}_{N,k}]^{T}$, can be obtained as
-$e_{k}$ $=(I_{N}\otimes(A-KCA))e_{k-1}$ (58)
-$~{}~{}~{}~{}+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA)e_{k-1}$
+$e_{k}$ $=(I_{N}\otimes(A-KCA))e_{k-1}+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA)e_{k-1}$ (34)
 $~~~~-(I_{N}\otimes K)G^{l}(1_{N}\otimes(C\omega_{k-1}+\nu_{k}))$
@@
 $~~~~-(I_{N}\otimes(I-KC)-(I_{N}\otimes K)$
-×$~{}~{}~{}~{}\times G^{l}(I_{N}\otimes C))(1_{N}\otimes\omega_{k-1})$
+$~~~~\times G^{l}(I_{N}\otimes C))(1_{N}\otimes\omega_{k-1})$
 $~~~~+(I_{N}\otimes K)(I_{Nr}-G^{l})(1_{N}\otimes\nu_{k}).$
-By denoting the matrices as (24), (25), and (26), $e_{k}$ has the following form
-$e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}).$
-Finally, $P_{k|k}$ can be calculated as (22) and (23) according to $P_{k|k}=E\{e_{k}e^{T}_{k}\}$
-Appendix D PROOF of Lemma 6
-Under the condition $0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}}$ and the fact $l_{ij}\leq 0,i\neq j$, it holds
-$1-\mu_{il}(l_{ii}+a_{il})\geq 0$ (59)
-and
-$-\mu_{il}l_{ij}\geq 0.$ (60)
-By combining (27), (28), (59), and (60), it can be concluded that the matrix $G$ is nonnegative. By utilizing Lemma 3 and observing (29), it follows
-$\rho(G)$ $\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}G_{ij}$
-$\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}g_{[ij,l]}$
-$\leq 1-\mu_{il}a_{il}.$
-Hence, the conclusion is drawn that $\rho(G)\leq 1$.
-Subsequently, it will be demonstrated that the eigenvalues of $G$ do not include the value $1$. By substituting the value $1$ into the characteristic polynomial of $G$, one has
-$|I-G|=|\Lambda(\mathcal{L}\otimes I_{r}+B)|.$
+By denoting the matrices as (8), (9), and (10), $e_{k}$ has the following form $e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}).$
+Item 4): $P_{k|k}$ can be calculated as (11) and (12) according to $P_{k|k}=E\{e_{k}e^{T}_{k}\}$ .
+Appendix B PROOF of Lemma 5
+Consider the row sum of the block matrix, and it holds
+$~~~\sum^{n}_{j=1}g_{[ij,l]}$ $=\Big(1-\mu_{il}a_{il}\Big)\otimes I_{r_{l}}.$ (35)
+Under the condition $0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}}$ and the fact $l_{ij}\leq 0,i\neq j$, it holds $1-\mu_{il}(l_{ii}+a_{il})\geq 0$ and $\mu_{il}l_{ij}\geq 0.$ By combining (13) and (14), it can be concluded that the matrix $G$ is nonnegative.
+By utilizing Lemma 3 and observing (35), it follows $\rho(G)\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}G_{ij}\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}g_{[ij,l]}\leq 1-\mu_{il}a_{il}.$ Hence, the conclusion is drawn that $\rho(G)\leq 1$.
+Subsequently, it will be demonstrated that the eigenvalues of $G$ do not include the value $1$.
+By substituting the value $1$ into the characteristic polynomial of $G$, one has $|I-G|=|\Lambda(\mathcal{L}\otimes I_{r}+B)|.$
 According to Definition 1, $L_{B}=\mathcal{L}\otimes I_{r}+B$ is irreducibly diagonally dominant.
@@
 In conclusion, it can be inferred that $\rho(G)<1$.
-Appendix E PROOF of Theorem 1
+Appendix C PROOF of Theorem 1
+By employing Lemma 2 and considering (8), it is evident that $\rho(\mathcal{A}(l))<1$ can guarantee the convergence of the distributed filter.
+The following results provide a lower bound of the fusion step number $l$.
 Since $\|G\|_{2}<1$, one can conclude that $G^{l}$ is uniformly upper-bounded for any positive $l$.
@@
 If $\mathcal{A}(l)$ is Schur stable, it can be concluded that $P_{k|k}$ is uniformly upper-bounded.
-Given Assumption 2, it is well known that the matrix $A-KCA$ is Schur stable [2].
+Given Assumption 1, it is well known that the matrix AKCAA-KCA is Schur stable [2].
 Applying Lemma 2, as $l$ tends to infinity, $\lim_{l\to\infty}G^{l}\to 0$.
@@
 In matrix theory, it is established that the spectral radius of $\rho(\mathcal{A}(l))$ is bounded by the spectral norm $\|\mathcal{A}(l)\|_{2}$, i.e., $\rho(\mathcal{A}(l))\leq\|\mathcal{A}(l)\|_{2}$.
-Using the identity $\|J_{1}\otimes J_{2}\|_{2}=\|J_{1}\|_{2}\|J_{2}\|_{2}$, it follows
-$\|\mathcal{A}(l)\|_{2}\leq\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}.$ (61)
-Let $\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}<1$. After some algebraic manipulations, one has
-$\|G\|^{l}_{2}<\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}.$ (62)
-Since $10<\|G\|_{2}<$ 1, there exists
-$l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}},$ (63)
-such that when $l>l_{0}$, $\mathcal{A}(l)$ is Schur stable. Consequently, it can be proven that $P_{k|k}$ is uniformly upper-bounded. Since
+Using the identity $\|J_{1}\otimes J_{2}\|_{2}=\|J_{1}\|_{2}\|J_{2}\|_{2}$, it follows $\|\mathcal{A}(l)\|_{2}\leq\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}.$ Let $\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}<1$.
+After some algebraic manipulations, one has
+$\|G\|^{l}_{2}<\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}.$
+Since $0<\|G\|_{2}<1$, there exists $l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}},$ such that when $l>l_{0}$, $\mathcal{A}(l)$ is Schur stable.
+Consequently, it can be proven that $P_{k|k}$ is uniformly upper-bounded.
+Since
 $P_{i,k|k}$ $=[0_{n\times n(i-1)},I_{n},0_{n\times n(N-i)}]P_{k|k}$ (36)
-×$~{}~{}~{}~{}\times[0^{T}_{n\times n(i-1)},I^{T}_{n},0^{T}_{n\times n(N-i)}]^{T},$
+$~~~~\times[0^{T}_{n\times n(i-1)},I^{T}_{n},0^{T}_{n\times n(N-i)}]^{T},$
 it can be found that $P_{i,k|k}$ is also uniformly upper-bounded.
-Appendix F PROOF of Theorem 2
-Under Assumption 1, 2, and 3, it can be concluded that $\mathcal{A}(l)$ and $\mathcal{A}_{cc}$ are Schur stable.
+Appendix D PROOF of Theorem 2
+Under Assumptions 1, 2, and 3, it can be concluded that $\mathcal{A}(l)$ and $\mathcal{A}_{cc}$ are Schur stable.
 In addition, $\Phi(l)$ and $\Phi_{cc}$ are also uniformly upper-bounded due to the boundedness of $K$, $C$, $Q$, $R$, and $G^{l}$.
 By utilizing Theorem 1 in [6], it can be proven that $P_{k|k}$ and $P_{cc,k|k}$ converge to $P^{(l)}$ and $P_{cc}$, respectively.
-Appendix G PROOF of Lemma 7
+Appendix E PROOF of Lemma 6
 First, consider the term $\mathcal{A}(l)P^{(l)}\mathcal{A}^{T}(l)$ in $P^{(l)}$, and one has
@@
 Performing the iteration (38) for $k$ times, it follows
-$P^{(l)}-P_{cc}$ $=\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}$ (68)
-$~{}~{}~{}~{}+\sum^{k-1}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$
-By performing an infinite number of iterations and utilizing the fact that $\mathcal{A}_{cc}$ is Schur stable, it holds
-$\lim_{k\to\infty}\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}=0$
-and
-$P^{(l)}-P_{cc}=\sum^{\infty}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$
-Next, similarly to (65), $\bar{\Phi}(l)$ in (67) can be calculated as (44).
-Appendix H PROOF of Theorem 3
+$P^{(l)}-P_{cc}$ $=\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}+\sum^{k-1}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$ (40)
+By performing an infinite number of iterations and utilizing the fact that $\mathcal{A}_{cc}$ is Schur stable, it holds $\lim_{k\to\infty}\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}=0$ and $P^{(l)}-P_{cc}=\sum^{\infty}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$ Next, similarly to (37), $\bar{\Phi}(l)$ in (39) can be calculated as (28).
+Appendix F PROOF of Theorem 3
 By calculating the spectral norm of (27), one has
@@
 $\leq M_{6}l^{Nr}\rho(G)^{l-Nr}.$
-By combining (69), (71), (74), and (75), a positive number $M_{1}$ can be designed such that
-$\|P^{(l)}-P_{cc}\|_{2}\leq M_{1}l^{Nr}\rho(G)^{l-Nr}.$ (76)
+By combining (41), (43), (46), and (47), a positive number $M_{1}$ can be designed such that $\|P^{(l)}-P_{cc}\|_{2}\leq M_{1}l^{Nr}\rho(G)^{l-Nr}.$
 Item 2): If $\|G\|_{2}<1$, (43) can be calculated as
 $\|\bar{\Phi}(l)\|_{2}\leq M_{3}\|G\|^{l}_{2}.$ (48)
-By combining (69), (74), and (77), a positive number $M_{2}$ can be obtained such that
-$\|P^{(l)}-P_{cc}\|_{2}\leq M_{2}\|G\|^{l}_{2}.$ (78)
-Appendix I PROOF of Lemma 8
-Based on Assumption 3 and Lemma 2, as $l$ tends to infinity, one has $\lim_{k\to\infty}G^{l}=0$. By combining (24), (25), and (26) and comparing them to (38), (39), and (40), this lemma can be proven.
-Appendix J PROOF of Theorem 4
-Utilizing Lemma 8, (22), and (36), it follows
-$\lim_{l\to+\infty}P_{k|k}=P_{cc,k|k}.$ (79)
-Based on (64), this conclusion can be drawn.
+By combining (41), (46), and (48), a positive number $M_{2}$ can be obtained such that $\|P^{(l)}-P_{cc}\|_{2}\leq M_{2}\|G\|^{l}_{2}.$